### v1

# Gradient-Free Adaptive Global Pruning for Pre-trained Language Models

Abstract
The transformative impact of large language models (LLMs) like LLaMA and GPT on natural language processing is countered by their prohibitive computational demands. Pruning has emerged as a pivotal compression strategy, introducing sparsity to enhance both memory and computational efficiency. Yet, traditional global pruning is impractical for LLMs due to scalability issues, while local pruning, despite its efficiency, leads to suboptimal solutions. Addressing these challenges, we propose Adaptive Global Pruning (AdaGP), a novel framework that redefines the global pruning process into manageable, coordinated subproblems, allowing for resource-efficient optimization with global optimality. AdaGP’s approach, which conceptualizes LLMs as a chain of modular functions and leverages auxiliary variables for problem decomposition, not only facilitates a pragmatic application on LLMs but also demonstrates significant performance improvements, particularly in high-sparsity regimes where it surpasses current state-of-the-art methods. The source code of AdaGP is available at https://github.com/BaiTheBest/AdaGP.
1 Introduction
Large language models Touvron et al. (2023); OpenAI (2023) have recently transformed the field of natural language processing (NLP) by delivering exceptional results across a variety of intricate language benchmarks Wei et al. (2022); Bommarito II and Katz (2022); Bubeck et al. (2023). Nonetheless, these models, with billions of parameters, generally necessitate significant computational resources. To make large language models (LLMs) more accessible, extensive efforts have been devoted to model compression of LLMs Xu and McAuley (2023); Ling et al. (2023); Bai et al. (2024), including pruning, quantization, knowledge distillation, and low-rank factorization, etc. Pruning, by introducing sparsity, jointly enhances memory and computational efficiency and offers unparalleled flexibility, seamlessly integrating with any LLMs, thus standing out as a highly effective and widely adopted compression strategy.
Figure 1: AdaGP decomposes the global pruning of LLMs into manageable subproblems by leveraging the chain of modules and auxiliary variables.
Model pruning has a long history LeCun et al. (1989) and has proven effective in applications related to vision and smaller language models Hoefler et al. (2021). However, conventional pruning techniques, which rely on global pruning and require loading the entire model into the same GPU Mallya and Lazebnik (2018); Singh and Alistarh (2020), become impractical for today’s LLMs due to their vast size. Recently, several local pruning methods, which consider compressing each layer separately, and the overall compressed model is then obtained by “stitching together” the individually compressed layers, are proposed for billion-scale LLMs. Frantar and Alistarh (2023) developed SparseGPT, an efficient unstructured pruning method for LLMs with hundreds of billions of parameters, achieving up to 60% parameter reduction with minimal performance loss. Sun et al. (2023) introduced a novel pruning criterion in Wanda, which evaluates weights by considering both magnitude and related input activations. However, despite its efficiency gains, local pruning only aims to minimize the local error for each specific layer under sparsity constraints and hence the final solution is suboptimal to the final goal of the global pruning. In other words, beyond the genuine goal of pruning which is just to align the input and output of the pruned and original models, local pruning over-aligns all the intermediate layers’ activations between them, resulting in a suboptimal solution, especially in high-sparsity regimes Singh and Alistarh (2020); Sung et al. (2023).
To address these challenges, namely, to achieve global pruning with low memory consumption, we propose Adaptive Global Pruning (AdaGP), that decomposes the global pruning objective into many subproblems, each of which can be solved using low resources and can coordinate each other toward the global pruning objective. More specifically, we first formulate LLMs as a composite function where the output of one module is the input of the next, based on which we reformulate the global pruning goal into its equivalent form, with auxiliary variables that facilitate its decomposition and its subproblems’ coordination. Then we propose an algorithm that achieves the alternating optimization of the subproblems of all the variables and auxiliary variables with low resource and global optimality, thanks to the closed-form solution of each subproblem. Empirically, we find that AdaGP can consistently improve the performance of local pruning methods, particularly in high sparsity regimes (> 60%), where the perplexity can be significantly decreased by up to around 80% as compared to the state-of-the-art local pruning methods, SparseGPT.
Furthermore, our AdaGP framework can be readily applicable to enhance the performance of most existing local pruning solvers, such as SparseGPT and Wanda, with marginal additional computational overhead. This adaptability ensures that our framework can be seamlessly integrated into a wide range of LLMs and pruning methods, making it a versatile tool and useful baseline for future research exploiting the sparsity of LLMs.
2 Related Work
Pruning, a pivotal concept in machine learning that introduces sparsity into neural networks, dates back to the 1980s LeCun et al. (1989). It gained renewed attention in the late 2010s, especially for deep neural networks, under the banner of reducing inference costs Han et al. (2015). LLM pruning techniques can broadly be categorized into structured and unstructured prunings.
Unstructured pruning looks at simplifying the complexity of LLMs by removing certain parameters regardless of the model’s inherent structure. This approach typically involves setting a threshold to nullify parameters below it, leading to a model with a non-uniform sparse structure. Frantar and Alistarh (2023) developed SparseGPT, an efficient unstructured pruning method for LLMs with hundreds of billions of parameters, achieving up to 60% parameter reduction with minimal performance loss. Sun et al. (2023) introduced a novel pruning criterion in Wanda, which evaluates weights by considering both magnitude and related input activations. This approach is beneficial in linear layers of LLMs, helping to identify and remove less significant weights. Tuli and Jha (2023) proposes DynaTran, a dynamic inference scheme for pruning activations at runtime, supported by a specially designed ASIC architecture, AccelTran, to enhance transformer inference throughput.
On the other hand, structured pruning involves the selective removal of groups of weights, where ‘group’ might mean blocks of weights, filters, attention heads, or other structures conducive to hardware acceleration. Ma et al. (2023) introduced the LLM-Pruner, a framework designed for structured pruning of LLMs, which utilizes a combination of first-order data and Hessian information for effective importance estimation. This aids in identifying crucial groups for pruning. Li et al. (2023) proposed LoSparse, a novel approach combining low-rank and sparse matrix approximations to balance pruning and expressive power. Tao et al. (2023) extended this concept to pruning hidden dimensions in LLMs, including embedding layers and attention heads. Kurtic et al. (2023) presented ZipLM, a structured compression method for LLMs, optimizing for compression and accuracy while considering specific hardware constraints.
Our work falls in the category of unstructured pruning of LLMs, where existing methods such as SparseGPT and Wanda only consider an entirely local pruning algorithm and suffer from sub-optimal performance. We discuss the limitations and challenges of entirely local pruning in Sec. 3.
3 Background
3.1 Global Pruning
Given a pre-trained neural network $f$ and inputs $\mathbf{X}$, global pruning aims to find a global sparsity mask $\mathbf{M}$ and possibly updated weights $\widehat{\mathbf{W}}$ to minimize the global loss between the final outputs of the uncompressed and compressed model:
${\min}_{\mathbf{M},\widehat{\mathbf{{W}}}}\;\;\mathcal{L}\big{(}f(\mathbf{X};\mathbf{M}\odot\widehat{\mathbf{W}}),f(\mathbf{X};\mathbf{W})\big{)},$ (1)
where $\odot$ denotes the element-wise multiplication. In addition to NP-hardness Blumensath and Davies (2008), however, a critical challenge in solving Eq. 1 for modern billion-scale LLMs is the huge memory requirement which prohibits storing the entire model on a single GPU. Although there exists toolkits, e.g., Paszke et al. (2019); Smith et al. (2022), that enable parallelism across multiple GPUs, such a setup may require resources that are not easily accessible, raising interests in developing other pruning methods that are functional regardless of the availability of computational resources.
3.2 Local Pruning
Local pruning circumvents the memory issue mentioned above by dividing the full model compression into sub-problems for each layer, $\ell$, and constructing a local loss to measure the $\ell_{2}$-error between the outputs of the uncompressed and compressed layers. Hence, the local pruning can be formulated by
$\text{min}_{\mathbf{M}_{\ell},\widehat{\mathbf{W}}_{\ell}}\|\mathbf{W}_{\ell}\cdot\mathbf{X}_{\ell}-(\mathbf{M}_{\ell}\odot\widehat{\mathbf{W}}_{\ell})\cdot\mathbf{X}_{\ell}\|_{2}^{2}.$ (2)
Although smaller than the global pruning, the local pruning still needs to optimize both the mask $\mathbf{M}_{\ell}$ and the remaining weights $\widehat{\mathbf{W}}_{\ell}$ and thus remains NP-hard. Therefore, exactly solving it for larger layers is unrealistic, leading all existing methods to resort to approximations.
Mask Selection & Weight Reconstruction.
A particularly popular approach is to separate the problem into mask selection and weight reconstruction Hubara et al. (2021); Kwon et al. (2022). Concretely, this means first choosing a pruning mask $\mathbf{M}$ according to some salient criterion, like the weight magnitude Zhu and Gupta (2017), and then optimizing the remaining un-pruned weights while keeping the mask unchanged. Importantly, once the mask is fixed, Eq. 2 turns into a linear regression problem that can be easily optimized.
Figure 2: Illustration of our adaptive-global pruning method compared to conventional global pruning and local pruning. We consider a two-layer neural network as an abstraction for simplicity. Global pruning (left) is memory prohibitive due to poor scalability. Local pruning (mid) considers pruning each layer independently, while inevitably sacrificing performance due to the ignorance of global supervision. Our adaptive global pruning (right) achieves global pruning with low memory cost by leveraging auxiliary variables and soft constraints.
Existing Solvers.
Early work Kingdon and Kingdon (1997) applied iterated linear regression to small networks. Recently, the AdaPrune approach Hubara et al. (2021) has shown good results for this problem on modern models via magnitude-based weight selection, followed by applying SGD steps to reconstruct the remaining weights. Follow-up works demonstrate that pruning accuracy can be further improved by removing the strict separation between mask selection and weight reconstruction. In addition, as we mentioned in Section 2, the work Frantar and Alistarh (2023); Sun et al. (2023) further advance in the scalability and pruning criterion of the pruning process.
3.3 What’s Wrong with Local Pruning?
As shown in Eq. 2, local pruning focuses on minimizing the error for each specific layer $\ell$ subject to sparsity constraints which are embedded in the mask $\mathbf{W}_{\ell}$ and weights $\widehat{\mathbf{W}}_{\ell}$. While the primary goal of the pruning process is to ensure that the input and output of the pruned model align closely with those of the original models, local pruning overly constrains the alignment of input and output of all the intermediate layers between the two models, leading to a performance degradation. This results in a solution that is suboptimal with respect to the global pruning problem.
4 Adaptive Global Pruning for LLMs
We present our proposed method AdaGP that can address the drawbacks of existing pruning methods by achieving a global pruning with low memory consumption. AdaGP decomposes the global pruning into many subproblems, each of which can be solved using low resources and can coordinate each other toward the global pruning objective.
4.1 Motivation
The development of AdaGP is motivated by the observation: LLMs can be formulated as a composite function such that the output of one module is the input of the next. This allows us to reformulate the global pruning goal into its equivalent form with auxiliary variables that enable the decomposition of the original problem into multiple subproblems, as detailed in Figure 2 and Section 4.2. Then we develop a resource-efficient algorithm that achieves global optimality via alternating optimization for each of the subproblems in which we obtain closed-form solutions for each of the steps. This has been illustrated in Figure 2 and Section 4.3.
4.2 A Unified Formulation of Pruning
In this section, we present the reformulation of the global pruning problem into an equivalent one by introducing auxiliary variables. This reformulation enables us to swiftly shift between two extremes of global pruning and local pruning and the decomposition of the problem into many independent subproblems.
The key idea behind our reformulation is to decouple the densely parametric parts (linear layers) from the non-parametric parts (activation function, self-attention, and layer norm, etc) using a splitting technique. Formally, given an LLM formulated as a composite function with dense layers $\ell\in[L]:=\{1,2,\ldots,L\}$, where $L$ is the total number of dense layers and define $\Omega\subseteq[L-1]$ to be a subset of layers for which we control the extent of globality. Note that there can be non-parameteric and potentially nonlinear layers between layers $\ell$ and $\ell+1$ in ${L}$ and we denote those layers by function $\phi_{\ell}(\cdot)$. In the reformulation, rather than feeding the output of a dense layer $\ell\in[L]$ directly into a non-parametric and potentially non-linear layer after it, we store the output of layer $\ell$ in a new variable $\mathbf{z}_{\ell}=\mathbf{W}_{\ell}\boldsymbol{a}_{\ell-1}$  [^1] For the sake of simplicity and clearer presentation, the bias term is omitted in the following equations where its exclusion does not lead to confusion., where $\boldsymbol{a}_{\ell-1}$ is the input to layer $\ell$ and can be represented by $\boldsymbol{a}_{\ell-1}=\phi_{\ell-1}(\mathbf{z}_{\ell-1})$. Note that by definition, $\boldsymbol{a}_{0}$ is the input data. Consequently, we can pose the pruning problem as:
$\begin{split}&\min_{\{\widehat{\mathbf{{W}}}_{\ell}\},\{\mathbf{M}_{\ell}\},\{\boldsymbol{a}_{\ell}\},\{\mathbf{z}_{\ell}\}}\mathcal{L}(\mathbf{z}_{L},\mathbf{y}),\\ &\text{s.t. \;\;}\mathbf{z}_{\ell}=(\mathbf{M}_{\ell}\odot\widehat{\mathbf{W}}_{\ell})\boldsymbol{a}_{\ell-1},\;\forall\;\ell\in[L],\\ &\quad\quad\boldsymbol{a}_{\ell}=\phi_{\ell}(\mathbf{z}_{\ell}),\;\forall\;\ell\in\Omega,\\ &\quad\quad\boldsymbol{a}_{\ell},\mathbf{z}_{\ell}=\boldsymbol{a}_{\ell}^{pre},\mathbf{z}_{\ell}^{pre},\;\forall\;\ell\in[L-1]\backslash\Omega,\end{split}$ (3)
where $[L-1]\backslash\Omega$ denotes the complement set of $\Omega$ and $\boldsymbol{a}_{\ell}^{pre}$ and $\mathbf{z}_{\ell}^{pre}$ denote the corresponding values of $\boldsymbol{a}_{\ell}$ and $\mathbf{z}_{\ell}$ obtained by running a forward-propagation from the pre-trained model respectively.
In our proposed reformulation above, the key to understanding its unified nature lies in the interpretation of the set $\Omega$. The versatility of our approach is illustrated in the following remark:
Remark 4.1.
For the given LLM and definition of $\Omega$, our formulation can seamlessly treat both global and local pruning as special cases depending on the choices of $\Omega$. Specifically:
  • When $\Omega=[L-1]$, solving our pruning formulation is equivalent to global pruning, accounting for inter-layer dependencies across the entire network.
  • When $\Omega=\emptyset$, the formulation simplifies to local pruning, considering each layer independently and ignoring inter-layer dependencies.
The ability to shift between these two extremes, and potentially any intermediate configuration, demonstrates the flexibility and comprehensiveness of our reformulation. This flexibility not only caters to a wide range of pruning strategies but also provides a unified framework to compare and contrast the effectiveness of different pruning methods under a consistent mathematical lens.
4.3 Algorithm Design
In this section, we introduce a choice of $\Omega$ that achieves a good trade-off between global and local pruning and propose an algorithm to solve the resulting problem. Specifically, we propose an adaptive approach where global pruning is applied within each decoder layer, while still adhering to a “local pruning” strategy between different decoder layers. Moreover, given that the feed-forward network (FFN) module in each decoder layer accounts for more than two-thirds of the total parameters Liu et al. (2024), our approach prioritizes the global pruning of these FFN layers. This strategy strikes a balance between the computational feasibility of pruning large-scale models and the effectiveness of the pruning process.
Formally, we select $\Omega$ to be the FFN layers, denoted by $\Omega_{\text{FFN}}$. As a result, we obtain the problem of interest by substituting $\Omega$ with $\Omega_{\text{FFN}}$ in Eq. 3. We have the following problem now:
$\begin{split}&\min_{\{\widehat{\mathbf{{W}}}_{\ell}\},\{\mathbf{M}_{\ell}\},\{\boldsymbol{a}_{\ell}\},\{\mathbf{z}_{\ell}\}}\mathcal{L}(\mathbf{z}_{L},\mathbf{y}),\\ &\text{s.t. \;}\mathbf{z}_{\ell}=(\mathbf{M}_{\ell}\odot\widehat{\mathbf{W}}_{\ell})\boldsymbol{a}_{\ell-1},\;\forall\;\ell\in[L],\\ &\quad\;\;\,\boldsymbol{a}_{\ell}=\phi_{\ell}(\mathbf{z}_{\ell}),\;\forall\;\ell\in\Omega_{\text{FFN}},\\ &\quad\;\boldsymbol{a}_{\ell},\mathbf{z}_{\ell}=\boldsymbol{a}_{\ell}^{pre},\mathbf{z}_{\ell}^{pre},\;\forall\;\ell\in[L-1]\backslash\Omega_{\text{FFN}},\end{split}$ (4)
Instead of solving Eq. 4 directly, we relax the constraints by adding an $\ell_{2}$-penalty function to the objective to obtain an unconstrained problem. In addition, by considering the “local pruning" between the decoder layer, we can simplify Eq. 4 to obtain a unconstrained subproblem for each of the decoder layer as:
$\begin{split}&\min_{\{\widehat{\mathbf{{W}}}_{\ell}\},\{\mathbf{M}_{\ell}\},\boldsymbol{a}_{\ell},\mathbf{z}_{\ell}}\alpha\|\mathbf{z}_{\ell+1}^{pre}-(\mathbf{M}_{\ell+1}\odot\widehat{\mathbf{W}}_{\ell+1})\boldsymbol{a}_{\ell}\|_{2}^{2}\\ &\quad\quad\quad+\beta\|\boldsymbol{a}_{\ell}-\phi_{\ell}(\mathbf{z}_{\ell})\|_{2}^{2},\\ &\quad\quad\quad+\alpha\|\mathbf{z}_{\ell}-(\mathbf{M}_{\ell}\odot\widehat{\mathbf{W}}_{\ell})\boldsymbol{a}_{\ell-1}^{pre}\|_{2}^{2},\end{split}$ (5)
where $\alpha$, $\beta$ are hyperparameters for the penalty of corresponding constraint in Eq. 4 and layers $\ell$ and $\ell+1$ correspond to the up-scaling and down-scaling linear layers in FFN. Note that for the last decoder layer, we replace $\mathbf{z}_{\ell+1}^{pre}$ with $y$.
In this work, we consider the alternating method to optimize our Eq. 5, i.e., optimize each variable while keeping the rest fixed. The careful and elaborate design of our Eq. 5 allows us to derive a closed-form solution for each of the variables:
Pruning Weight.
First consider optimizing Eq. 5 with respect to $\mathbf{M}_{\ell}$ and $\widehat{\mathbf{W}}_{\ell}$. For each linear layer $\ell$ in a FFN module, the optimal solution minimizes $\|\mathbf{z}_{\ell}-(\mathbf{M}_{\ell}\odot\widehat{\mathbf{W}}_{\ell})\boldsymbol{a}_{\ell-1}\|_{2}^{2}$. To solve it, the first step is to write $\mathbf{z}_{\ell}$ as $\mathbf{W}_{\ell}\boldsymbol{a}_{\ell-1}$, where $\mathbf{W}_{\ell}=\mathbf{z}_{\ell}\boldsymbol{a}_{\ell-1}^{\dagger}$ ($\dagger$ denotes the pseudo-inverse.). We get $\|\mathbf{W}_{\ell}\boldsymbol{a}_{\ell-1}-(\mathbf{M}_{\ell}\odot\widehat{\mathbf{W}}_{\ell})\boldsymbol{a}_{\ell-1}\|_{2}^{2}$ by plugging $\mathbf{z}_{\ell}$ back in original loss. This aligns with the pruning objective of Eq. 2 and can be analytically solved by existing pruning solver e.g., SparseGPT. The superscript of “$pre$” for $\boldsymbol{a}_{\ell-1}$ is omitted in this section for simpler notation.
Updating Activation.
Minimization for $\boldsymbol{a}_{\ell}$ is a simple least-squares problem similar to weight pruning. However, in this case, the matrix $\boldsymbol{a}_{\ell-1}$ appears in two penalty terms in Eq. 5, so we must minimize $\alpha\|\mathbf{z}_{\ell+1}^{pre}-(\mathbf{M}_{\ell+1}\odot\widehat{\mathbf{W}}_{\ell+1})\boldsymbol{a}_{\ell}\|_{2}^{2}+\beta\|\boldsymbol{a}_{\ell}-\phi_{\ell}(\mathbf{z}_{\ell})\|_{2}^{2}$ for $\boldsymbol{a}_{\ell}$, holding all other variables fixed. By following a very similar idea to Ridge regression, the new value of $\boldsymbol{a}_{\ell}$ is given by:
$(\alpha\mathbf{W}_{\ell+1}^{\intercal}\mathbf{W}_{\ell+1}+\beta\mathbf{I})^{-1}(\alpha\mathbf{W}_{\ell+1}^{\intercal}\mathbf{z}_{\ell+1}^{pre}+\beta\phi_{\ell}(\mathbf{z}_{\ell})),$ (6)
where $\mathbf{W}_{\ell}$ denotes the updated weight matrix after pruning, i.e., $\mathbf{W}_{\ell}\coloneqq\mathbf{M}_{\ell}\odot\widehat{\mathbf{W}}_{\ell}$.
Updating Output.
The update for $\mathbf{z}_{\ell}$ requires minimizing the following loss:
$\beta\|\boldsymbol{a}_{\ell}-\phi_{\ell}(\mathbf{z}_{\ell})\|_{2}^{2}+\alpha\|\mathbf{z}_{\ell}-(\mathbf{M}_{\ell}\odot\widehat{\mathbf{W}}_{\ell})\boldsymbol{a}_{\ell-1}^{pre}\|_{2}^{2}.$ (7)
This problem can be challenging because of the potentially non-linear function $\phi_{\ell}(\cdot)$. Fortunately, as the non-linearity $\phi_{\ell}(\cdot)$ between two consecutive FFN layers in the transformer is typically ReLU, which works entry-wise on its argument, the entries in $\mathbf{z}_{\ell}$ are decoupled. As a result, solving Eq. 7 is particularly easy when $\phi_{\ell}$ is piece-wise linear (as in the case of ReLU). It can be solved in closed-form following a simple if-then logic. Specifically, one only needs to compute two solutions as:
$\begin{split}&\mathbf{z}_{\ell}^{(1)}=(\mathbf{M}_{\ell}\odot\widehat{\mathbf{W}}_{\ell})\boldsymbol{a}_{\ell-1}^{pre},\\ &\mathbf{z}_{\ell}^{(2)}=(\alpha+\beta)^{-1}\cdot\big{(}\beta\boldsymbol{a}_{\ell}+\alpha\mathbf{z}_{\ell}^{(1)}\big{)},\end{split}$ (8)
where the first solution corresponds to those entries of $\mathbf{z}_{\ell}$ that are negative (reduced to zero by ReLU), and the second solution corresponds to those entries of $\mathbf{z}_{\ell}$ that are non-negative (i.e., $\phi_{\ell}(\mathbf{z}_{\ell})=\mathbf{z}_{\ell}$). For more sophisticated choices of $\phi_{\ell}$, including smooth sigmoid curves, the problem can be solved quickly with a lookup table of pre-computed solutions because each entry only depends on two inputs.
Remark 4.2 (Convergence to global optimum with AdaGP).
Consider the unconstrained problem given in Eq. 5, we observe that, if the activation function $\phi$ is ReLU, (1) the objective function is convex with respect to each variable when all others are fixed, and (2) closed-form solutions exist for each of the variable updates in the alternating optimization scheme. The proposed algorithm resembles a multi-block alternating direction method of multipliers (ADMM) that is shown to converge to the global optimum in many applications.
4.4 Time Complexity Analysis
The proposed AdaGP is composed of three main steps, with the overall time complexity being the sum of the complexities of the individual steps. Consider one epoch of AdaGP, in the step of weights pruning, the time complexity is dominated by the computation of the pseudo-inverse of matrix $\boldsymbol{a}_{\ell}$ with dimensions $n\times h$ (where $n$ is the number of samples and $h$ is the hidden dimension of the transformer) and thus given by $O(nh^{2})$, assuming $n<h$. The exact pruning step, when employing SparseGPT as the solver, has a complexity of $O(h^{3})$. Combining both sub-steps, weight pruning has a total time complexity of $O(h^{3})$. The second step, updating activation, is dominated by a matrix inversion as shown in Eq. 6. Given that the weight matrix $\mathbf{W}_{\ell}$ is of size $h\times h$, this inversion has a time complexity of $O(h^{3})$. The third step, updating outputs, does not involve matrix inversion and its complexity does not exceed that of the previous steps. Thus, the overall time complexity of the algorithm is bounded by $O(h^{3})$. Consequently, our method exhibits a per-epoch time complexity that is on par with SparseGPT.
5 Experiments
Experiments Setup.
We implement AdaGP in PyTorch Paszke et al. (2019) and use the HuggingFace Transformers library Wolf et al. (2019) for handling models and datasets. All pruning experiments are conducted on NVIDIA A100 GPUs. Similar to Frantar and Alistarh (2023); Sun et al. (2023), we sparsify each transformer decoder layer sequentially in order, which significantly reduces memory requirements, and only apply global pruning to FFN modules. For the calibration data, we follow Frantar and Alistarh (2023) and use 2048-token segments, randomly chosen from the first shard of the C4 Raffel et al. (2020) dataset. This represents generic text data crawled from the internet and ensures our experiments remain zero-shot as no task-specific data is seen during pruning.
Models, Datasets & Evaluation.
Following Frantar and Alistarh (2023), we primarily work with the OPT model family Zhang et al. (2022) to study scaling behavior at various model sizes, but our method can be easily applied to other LLMs such as Llama Touvron et al. (2023). While our focus lies on the very largest variants, we also show some results on smaller models to provide a broader picture. In terms of metrics, we mainly focus on perplexity, which is known to be a challenging and stable metric that is well-suited for evaluating the accuracy of compression methods Yao et al. (2022); Dettmers and Zettlemoyer (2023). We consider the test sets of raw-WikiText2 Merity et al. (2016) and PTB Marcus et al. (1994) as well as a subset of the C4 validation data, all popular benchmarks in LLM compression literature Yao et al. (2022); Park et al. (2022); Frantar and Alistarh (2023); Sun et al. (2023). For additional interpretability, we also provide zero-shot accuracy results following the same setup of Sun et al. (2023), which is based on the popular EleutherAI-eval harness Gao et al. (2023).
OPT-1.3b (WikiText2: 14.62; PTB: 20.29; C4: 16.07)
Sparsity
70% 80% 90% 2:4
Dataset
WikiText2
PTB
C4
WikiText2
PTB
C4
WikiText2
PTB
C4
WikiText2
PTB
C4
Magnitude
6420.80
4828.13
3435.99
9998.71
1.1e4
5347.89
8209.13
1.0e4
4917.02
96.68
133.92
48.08
Wanda
21.56
34.77
25.78
142.20
146.76
142.24
5692.65
4751.69
4501.73
15.63
24.04
18.23
SparseGPT
18.04
28.19
21.45
69.67
93.36
60.83
2596.70
2361.86
1363.08
15.11
23.71
17.88
AdaGP
17.82
27.72
20.99
58.92
85.33
58.36
1350.31
1192.36
655.76
14.97
23.40
17.67
OPT-2.7b (WikiText2: 12.47; PTB: 17.97; C4: 14.32)
Sparsity
70% 80% 90% 2:4
Dataset
WikiText2
PTB
C4
WikiText2
PTB
C4
WikiText2
PTB
C4
WikiText2
PTB
C4
Magnitude
1691.74
1237.08
1415.02
1.0e4
7916.69
6050.07
7.9e5
5.3e5
4.7e5
272.34
308.55
267.70
Wangda
88.61
140.09
90.06
6140.81
4746.96
5678.66
3.0e4
3.5e4
2.4e4
13.66
21.67
16.10
SparseGPT
13.79
21.18
16.18
24.32
37.82
25.92
2662.74
2285.01
1776.08
12.62
19.28
15.12
AdaGP
13.82
21.07
16.14
23.87
37.09
24.90
1200.12
759.11
527.70
12.63
19.31
15.12
OPT-6.7b (WikiText2: 10.86; PTB: 15.77; C4: 12.71)
Sparsity
70% 80% 90% 2:4
Dataset
WikiText2
PTB
C4
WikiText2
PTB
C4
WikiText2
PTB
C4
WikiText2
PTB
C4
Magnitude
7054.21
5437.44
4850.25
7937.49
5971.86
6031.54
2.4e4
2.5e4
2.1e4
64.11
92.23
82.67
Wanda
54.95
129.73
116.67
1493.58
1196.93
996.00
2.1e4
2.0e4
1.8e4
11.86
18.54
14.77
SparseGPT
12.27
18.90
15.28
31.04
51.26
29.42
8871.24
5713.57
3797.20
11.30
16.90
13.51
AdaGP
12.16
18.39
14.93
23.96
39.32
26.97
2095.85
1842.48
953.44
11.07
16.73
13.42
OPT-13b (WikiText2: 10.13; PTB: 14.52; C4: 12.06)
Sparsity
70% 80% 85% 2:4
Dataset
WikiText2
PTB
C4
WikiText2
PTB
C4
WikiText2
PTB
C4
WikiText2
PTB
C4
Magnitude
9037.12
7734.58
5909.47
1.1e4
9140.88
6340.22
1.3e4
1.3e4
9087.50
67.07
110.77
52.61
Wanda
30.94
39.26
33.31
4216.04
2894.77
2450.57
1.1e4
1.1e4
7244.96
10.33
15.35
12.54
SparseGPT
10.89
16.35
13.39
21.42
33.62
21.01
8408.03
6380.30
3416.23
10.20
15.14
12.48
AdaGP
10.96
16.57
13.38
19.07
28.77
19.29
2052.27
1536.51
538.61
10.21
15.15
12.48
OPT-30b (WikiText2: 9.56; PTB: 14.04; C4: 11.45)
Sparsity
70% 80% 90% 2:4
Dataset
WikiText2
PTB
C4
WikiText2
PTB
C4
WikiText2
PTB
C4
WikiText2
PTB
C4
Magnitude
8691.40
4769.89
4732.66
8941.81
5292.98
5092.26
3.8e7
3.0e7
1.4e7
517.71
374.54
991.95
Wanda
7766.61
5547.45
5741.74
8770.33
6020.70
7132.20
6354.15
4296.37
4654.27
10.94
16.64
13.50
SparseGPT
9.58
14.41
11.93
16.49
22.01
17.67
5747.87
5169.50
3555.24
9.59
14.05
11.60
AdaGP
9.56
14.40
11.94
15.61
19.64
16.61
3050.63
2712.39
1758.63
9.57
14.04
11.60
OPT-66b (WikiText2: 9.34; PTB: 13.36; C4: 10.99)
Sparsity
70% 80% 90% 2:4
Dataset
WikiText2
PTB
C4
WikiText2
PTB
C4
WikiText2
PTB
C4
WikiText2
PTB
C4
Magnitude
OOM
OOM
OOM
OOM
OOM
OOM
OOM
OOM
OOM
OOM
OOM
OOM
Wanda
OOM
OOM
OOM
OOM
OOM
OOM
OOM
OOM
OOM
OOM
OOM
OOM
SparseGPT
9.45
13.64
11.37
28.27
57.41
26.26
7803.10
6594.88
4433.35
9.38
13.37
11.12
AdaGP
9.37
13.66
11.37
16.45
21.00
17.70
7504.17
5644.65
3683.91
9.36
13.36
11.12
Table 1: Perplexity of pruned OPT models in high sparsity regimes (> 60%); the lower the perplexity, the better.
OPT-6.7b
Sparsity
Method
BoolQ
RTE
HellaSwag
WinoGrande
ARC-e
ARC-c
OBQA
Mean
Dense
66.12
56.03
50.49
65.27
65.72
30.63
27.60
51.69
70%
SparseGPT
61.74
54.87
48.46
63.85
64.31
29.27
25.40
49.70
AdaGP
60.61
54.51
48.8
62.9
64.14
30.03
26.60
49.66
80%
SparseGPT
55.08
48.38
42.22
59.43
57.79
25.85
21.40
44.31
AdaGP
58.69
51.26
43.78
59.67
58.38
26.88
22.00
45.81
90%
SparseGPT
38.53
53.07
26.00
48.07
26.81
21.67
14.40
32.65
AdaGP
46.48
52.71
26.21
51.70
27.44
19.71
13.40
33.95
2:4
SparseGPT
62.11
54.87
49.58
64.4
65.32
30.63
27.60
50.64
AdaGP
63.82
55.96
49.55
64.56
65.70
30.55
27.60
51.11
OPT-13b
Sparsity
Method
BoolQ
RTE
HellaSwag
WinoGrande
ARC-e
ARC-c
OBQA
Mean
Dense
65.87
57.76
52.44
66.02
67.82
33.46
28.62
53.14
70%
SparseGPT
63.03
54.87
50.89
65.43
67.47
32.85
26.40
51.56
AdaGP
63.85
55.23
50.73
65.67
66.46
31.83
27.20
51.57
80%
SparseGPT
59.72
52.35
46.82
61.48
62.50
31.23
21.80
47.99
AdaGP
60.89
53.07
46.19
62.12
62.21
30.38
23.00
48.27
90%
SparseGPT
47.49
52.71
33.17
51.54
39.98
21.33
17.80
37.72
AdaGP
53.43
52.71
38.19
52.96
46.68
25.26
17.40
40.95
2:4
SparseGPT
64.71
54.51
51.92
65.98
67.47
32.85
28.20
52.23
AdaGP
65.63
55.23
51.93
65.27
67.8
33.45
28.40
52.53
OPT-30b
Sparsity
Method
BoolQ
RTE
HellaSwag
WinoGrande
ARC-e
ARC-c
OBQA
Mean
Dense
70.46
61.82
54.27
69.02
70.47
35.49
30.20
55.96
70%
SparseGPT
68.78
58.48
53.83
67.64
69.15
34.30
29.60
54.54
AdaGP
69.11
61.73
53.97
68.43
69.78
34.73
29.80
55.36
80%
SparseGPT
64.86
60.65
49.73
61.40
61.91
31.74
24.20
50.64
AdaGP
65.41
59.57
50.65
61.96
62.71
32.25
26.50
51.29
90%
SparseGPT
37.83
53.79
25.96
49.88
26.47
20.22
12.60
32.39
AdaGP
43.55
52.35
26.32
50.04
27.31
20.56
14.00
33.45
2:4
SparseGPT
69.54
58.12
54.08
67.25
69.95
34.81
29.80
54.79
AdaGP
69.67
58.12
54.08
68.88
70.45
35.49
30.20
55.27
Table 2: Accuracy (%) of zero-shot tasks; the higher the accuracy, the better.
Comparison Methods.
We compare against the standard magnitude pruning baseline Zhu and Gupta (2017) applied locally. We also compare with two state-of-the-art local pruning methods, SparseGPT Frantar and Alistarh (2023) and Wanda Sun et al. (2023).
5.1 Results and Analyses
Pruning vs. Model Sizes.
We begin by exploring the pruning capabilities of AdaGP across various model sizes in comparison to the baseline methods. For each model, we consider unstructured sparsity ranging from 70% to 90% with a 10% increment, as well as a full 2:4 semi-structured sparsity. We conduct a sensitivity study on the calibration sample sizes (see Appendix A.4) and use calibration sample sizes between 32 and 64 for all experiments. Moreover, except for the OPT-125m model, we prune the first 50% of the transformer decoder layers in each model to achieve a balance between the computation resources and showcasing the performances. Detailed results can be found in Table 1 and Table 3 (see Appendix A.1). The perplexity results on the three datasets for the un-pruned dense models are reported next to the names of the models. Note that we exclude the results for OPT-350m and adjust 90% unstructured sparsity for OPT-1.3b to 85% unstructured sparsity for numerical stability. We provide a discussion on the selection of hyperparameters of $\alpha$ and $\beta$ (used in Eq. 5) in Appendix A.5.
From the table, it shows that perplexity increases with increasing sparsity. Moreover, we observe a trend of decreasing perplexity for SparseGPT and AdaGP at the same sparsity with increasing model sizes. However, such a trend is not obvious for Magnitude and Wanda. We also observe that SparseGPT and AdaGP consistently outperform Magnitude and Wanda by a significant margin. For smaller sparsity, AdaGP achieves comparable perplexity to SparseGPT. As we increase the sparsity, AdaGP starts to demonstrate noticeable improvements over SparseGPT. In numerous instances, AdaGP achieves perplexity reductions of more than 50% compared to SparseGPT.
Computation Time vs. Model Sizes.
Next, we study how the computation time for a single epoch of SparseGPT and AdaGP varies with different model sizes, as illustrated in Figure 3. We observe that the time taken for a single epoch increases with increasing model sizes. The rate at which the time taken increases is comparable for SparseGPT and AdaGP. Additionally, although SparseGPT executes only one epoch for each layer to prune, whereas AdaGP may execute multiple epochs depending on the parameter configuration, we find that in most cases, few epochs (less than 10) of AdaGP suffice. This suggests that the proposed AdaGP remains computationally efficient.
Figure 3: Time for one step of SparseGPT and AdaGP.
Zero-shot Experiments.
To conclude the evaluations and discussions, we show the results for several zero-shot tasks in Table 2 comparing SparseGPT and AdaGP and three model sizes. These evaluations are known to be relatively noisy Dettmers et al. (2022), but more interpretable. We also report the results for zero-shot tasks from the un-pruned dense models in the “Dense" row. We see that the accuracy of both methods decreases with increasing sparsity, which is expected, as more parameters are pruned. A similar trend of increasing accuracy with increasing model size is observed too. Across all the tasks, OBQA and ARC-c remain the most challenging ones as the accuracy for both methods remain around or below 30% while both methods perform well for BoolQ, RTE, WinoGrande, and ARC-e. In general, AdaGP is able to achieve higher accuracy in almost all tasks across the three models of different sizes.
6 Conclusion
Our work presents AdaGP, a cutting-edge framework poised to redefine the compression of LLMs through sparsity. By adeptly circumventing the scalability issues of global pruning and addressing the local suboptimality of existing methods, AdaGP stands as a significant advancement in the field. Our empirical results affirm its efficacy, particularly in high-sparsity environments, where it achieves a notable reduction in perplexity, thereby setting a new precedent for model compression. The versatility and minimal computational overhead of AdaGP complement its integration with current pruning technologies, underscoring its potential as a universal tool for enhancing the performance and accessibility of LLMs.
7 Limitations
While AdaGP marks a significant step forward in efficient pruning of large language models, it is important to acknowledge the inherent trade-off associated with any model compression technique. Firstly, while our method reduces the complexity of LLMs and enhances computational efficiency, there is an inevitable balance between sparsity and performance that requires careful calibration. Additionally, the effectiveness of AdaGP, similar to any pruning method, may vary across different models and tasks, and its generalizability to all scenarios remains an area for further exploration. Lastly, our approach assumes certain structural properties of the neural network, such as layer-wise decomposability, which may not hold for all architectures. These limitations, while not undermining the contributions of AdaGP, suggest avenues for future research and refinement.
Acknowledgement
This research used resources of the Argonne Leadership Computing Facility at Argonne National Laboratory, which is supported by the Office of Science of the U.S. Department of Energy under contract DE-AC02-06CH11357. We gratefully acknowledge the computing resources provided on Swing, a high-performance computing cluster operated by the Laboratory Computing Resource Center at Argonne National Laboratory.
Government license
The submitted manuscript has been created by UChicago Argonne, LLC, Operator of Argonne National Laboratory ("Argonne”). Argonne, a U.S. Department of Energy Office of Science laboratory, is operated under Contract No. DE-AC02-06CH11357. The U.S. Government retains for itself, and others acting on its behalf, a paid-up nonexclusive, irrevocable worldwide license in said article to reproduce, prepare derivative works, distribute copies to the public, and perform publicly and display publicly, by or on behalf of the Government. The Department of Energy will provide public access to these results of federally sponsored research in accordance with the DOE Public Access Plan (http://energy.gov/downloads/doe-public-access-plan)
Appendix A Appendix
A.1 OPT-125m
Detailed results for OPT-125m are in Table 3. For a thorough investigation, we additionally consider unstructured sparsity of 50% and 60% as well as pruning the full model (instead of the first 50% of the transformer decoder layers). Similar to other models, we report the perplexity results for the un-pruned dense model next to the name of the model in the table. In particular, we see that SparseGPT and AdaGP outperform Magnitude and Wanda with a significant margin across different level of sparsity. AdaGP achieves similar perplexity with SparseGPT at smaller sparsity, but demonstrates much better perplexity for larger sparsity. Similar perplexity trends are observed across the three datasets, although, PTB, having the highest perplexity for each sparsity and method, seems to be the most challenging dataset among the three.
OPT-125m (WikiText2: 27.66; PTB: 38.99; C4: 26.56)
Sparsity Method WikiText2 PTB C4 Sparsity Method WikiText2 PTB C4
50% Magnitude 193.39 276.33 141.03 80% Magnitude 4890.96 4121.49 3213.85
Wanda 39.58 57.44 35.12 Wanda 1912.45 2512.93 1066.86
SparseGPT 37.49 55.91 33.63 SparseGPT 2072.12 1952.85 1050.83
AdaGP 37.09 55.82 33.40 AdaGP 1358.10 1418.09 654.54
60% Magnitude 920.29 1146.49 598.45 90% Magnitude 6613.18 5380.80 4475.29
Wanda 76.30 106.92 61.33 Wanda 4940.89 4337.27 3126.02
SparseGPT 62.22 87.70 49.85 SparseGPT 6131.57 6963.27 2443.33
AdaGP 59.28 84.36 47.78 AdaGP 5291.64 5067.41 2003.09
70% Magnitude 3806.96 3429.35 2263.37 2:4 Magnitude 342.51 798.99 223.19
Wanda 351.83 412.52 248.94 Wanda 83.81 113.01 65.78
SparseGPT 239.26 265.83 156.33 SparseGPT 61.44 92.30 51.42
AdaGP 208.46 255.75 137.72 AdaGP 56.55 81.46 46.14
Table 3: Perplexity on pruned OPT-125m; the lower the perplexity, the better
A.2 Pseudo-code of AdaGP
Input: Pre-trained weight matrices $\mathbf{W}_{\ell}$ and $\mathbf{W}_{\ell+1}$, inputs $\boldsymbol{a}_{\ell-1}^{pre}$ and $\boldsymbol{a}_{\ell}^{pre}$, and outputs $\mathbf{z}_{\ell}^{pre}$ and $\mathbf{z}_{\ell+1}^{pre}$ for FFN’s up-scaling and down-scaling layers $\ell$ and $\ell+1$; target sparsity $\rho$; constraint weight hyperparameters $\alpha$, $\beta$.
1AdaGP():
2       Initialize $\mathbf{z}_{\ell}=\mathbf{z}_{\ell}^{pre}$, $\boldsymbol{a}_{\ell}=\boldsymbol{a}_{\ell}^{pre}$triangleright Initialize slack variables
3       Pre-compute and cache $\boldsymbol{a}_{\ell-1}^{\dagger}$ = pseudo-inverse($\boldsymbol{a}_{\ell-1}^{pre}$)
4       for step $i=1,\cdots,K$ do
5             $\mathbf{W}_{\ell}=\mathbf{z}_{\ell}\boldsymbol{a}_{\ell-1}^{\dagger}$, $\mathbf{W}_{\ell+1}=\mathbf{z}_{\ell+1}^{pre}\boldsymbol{a}_{\ell}^{\dagger}$
6             $\mathbf{M}_{\ell},\widehat{\mathbf{W}}_{\ell}=\operatorname*{arg\,min}\|\mathbf{W}_{\ell}\boldsymbol{a}_{\ell-1}^{pre}-(\mathbf{M}_{\ell}\odot\widehat{\mathbf{W}}_{\ell})\boldsymbol{a}_{\ell-1}^{pre}\|_{2}^{2}$triangleright Prune layer $\ell$ by SparseGPT solver
7             $\mathbf{M}_{\ell+1},\widehat{\mathbf{W}}_{\ell+1}=\operatorname*{arg\,min}\|\mathbf{W}_{\ell+1}\boldsymbol{a}_{\ell}-(\mathbf{M}_{\ell+1}\odot\widehat{\mathbf{W}}_{\ell+1})\boldsymbol{a}_{\ell}\|_{2}^{2}$triangleright Prune layer $\ell+1$ by SparseGPT solver
8             $\mathbf{W}_{\ell+1}=\mathbf{M}_{\ell+1}\odot\widehat{\mathbf{W}}_{\ell+1}$, $\mathbf{W}_{\ell}=\mathbf{M}_{\ell}\odot\widehat{\mathbf{W}}_{\ell}$
9             $\boldsymbol{a}_{\ell}=(\alpha\mathbf{W}_{\ell+1}^{\intercal}\mathbf{W}_{\ell+1}+\beta\mathbf{I})^{-1}(\alpha\mathbf{W}_{\ell+1}^{\intercal}\mathbf{z}_{\ell+1}^{pre}+\beta\phi_{\ell}(\mathbf{z}_{\ell}))$triangleright Update activations
10             $\mathbf{z}_{\ell}^{(1)}=\mathbf{W}_{\ell}\boldsymbol{a}_{\ell-1}^{pre},\;\mathbf{z}_{\ell}^{(2)}=(\alpha+\beta)^{-1}\cdot\big{(}\beta\boldsymbol{a}_{\ell}+\alpha\mathbf{z}_{\ell}^{(1)}\big{)},$
11             for $j=1,\cdots,n$ in parallel do
12                   if ${(\mathbf{z}_{\ell})}_{j}<0$ then
13                         ${(\mathbf{z}_{\ell})}_{j}={(\mathbf{z}_{\ell}^{(1)})}_{j}$triangleright Update outputs
14
15                  else
16                        ${(\mathbf{z}_{\ell})}_{j}={(\mathbf{z}_{\ell}^{(2)})}_{j}$triangleright Update outputs
17
18
19      return $\mathbf{W}_{\ell}$, $\mathbf{W}_{\ell+1}$
Algorithm 1 Adaptive Global Pruning for the FFN module of an LLM.
The Adaptive Global Pruning (AdaGP) algorithm presented in Algorithm 1 focuses on the optimization of the Feed-Forward Network (FFN) module within a Large Language Model (LLM). The key inputs to the algorithm include the pre-trained weight matrices, inputs and outputs for both the up-scaling and down-scaling linear layers of the FFN, along with a set of hyperparameters and a target sparsity. The goal of AdaGP is to achieve a targeted level of sparsity in the linear layers without significantly compromising the model’s performance.
Initiating with the pre-trained weights, AdaGP employs a series of pruning and activation update steps across $K$ epochs. In each epoch, it solves optimization problems to prune the weights in both up-scaling and down-scaling layers, followed by updating the activation variables. The utilization of SparseGPT solver for pruning and the strategic update of activations ensure that the pruned network behaves as similarly as possible to that of the original network. The final output of the algorithm is a pair of pruned weight matrices for the consecutive layers, which are expected to deliver comparable, if not improved, performance with a reduced number of non-zero parameters.
Figure 4: Visualization of the AdaGP Algorithm Applied to the FFN Module in a transformer decoder. This diagram illustrates the flow of data through the layers of the decoder, where AdaGP performs global pruning to enhance sparsity and computational efficiency without compromising the model’s performances. Each decoder layer is shown with its core components: Masked Multi-Head Attention and Feed Forward network, with AdaGP focusing on the latter for pruning. For other linear layers outside of the FFN modules, AdaGP follows the same local pruning as SparseGPT does to prune them.
A.3 Visualization of AdaGP for FFN Global Pruning
Figure 4 presents a schematic representation of the AdaGP algorithm applied to the Feed-Forward Network (FFN) module within a transformer model’s decoder layer. This visualization highlights the process of global pruning applied to the FFN module, aimed at reducing the model’s complexity while preserving its ability to capture essential patterns in the data.
The AdaGP method operates within the decoder block, selectively pruning the weights of the FFN module. Each decoder block comprises a series of operations. It starts with masked multi-head attention and follows by additive normalization. Subsequently, it passes through the feed-forward network in which the pruning happens with AdaGP. The figure illustrates the flow of information from an input token through successive layers of the decoder. AdaGP not only enhances computational efficiency but also has the potential to improve model generalization by eliminating redundant parameters.
A.4 Calibration Samples
Figure 5 presents how perplexity changes with the calibration sample sizes on the datasets PTB and C4, respectively. On one hand, in both plots, as the number of calibration samples increases, the perplexity decreases for both SparseGPT and AdaGP. This suggests that having more calibration samples can be beneficial in the pruning process. Perplexity decreases more rapidly from 8 samples to 32 samples. Beyond 32 samples, the rate at which perplexity decreases starts to slow down. For PTB dataset, the perplexity even increases from 64 samples to 128 samples. On the other hand, increasing the number of calibration samples requires more computational resources, e.g., memory and computation time, in the execution. A calibration sample size should be between 32 and 64 can therefore strike a balance between good performances and computational efficiency. Lastly, we want to point out that the figures show that AdaGP starts to achieve better perplexity than SparseGPT does from 32 samples and remains better for 64 and 128 samples.
Figure 5: Sensitivity of OPT-2.7b on the calibration sample sizes for datasets PTB and C4.
A.5 Hyperparameter $\alpha$ and $\beta$ Selection
Hyperparameters $\alpha$ and $\beta$ are used in Eq. 5. We select $\alpha$ and $\beta$ from the set $\{0.1,1.0,5.0\}$ and perform a study on model OPT-2.7b to understand the impact of the hyperparameters. We determine that $\alpha=\beta=0.1$ give the best and numerically stable performances.

### v3

@@
-Gradient-Free Adaptive Global Pruning for Pre-trained Language Models
+SparseLLM: Towards Global Pruning of Pre-trained Language Models
 Abstract
@@
 Yet, traditional global pruning is impractical for LLMs due to scalability issues, while local pruning, despite its efficiency, leads to suboptimal solutions.
-Addressing these challenges, we propose Adaptive Global Pruning (AdaGP), a novel framework that redefines the global pruning process into manageable, coordinated subproblems, allowing for resource-efficient optimization with global optimality.
-AdaGP’s approach, which conceptualizes LLMs as a chain of modular functions and leverages auxiliary variables for problem decomposition, not only facilitates a pragmatic application on LLMs but also demonstrates significant performance improvements, particularly in high-sparsity regimes where it surpasses current state-of-the-art methods.
-The source code of AdaGP is available at https://github.com/BaiTheBest/AdaGP.
+Addressing these challenges, we propose SparseLLM, a novel framework that redefines the global pruning process into manageable, coordinated subproblems, allowing for resource-efficient optimization with global optimality.
+SparseLLM’s approach, which conceptualizes LLMs as a chain of modular functions and leverages auxiliary variables for problem decomposition, not only facilitates a pragmatic application on LLMs but also demonstrates significant performance improvements, particularly in high-sparsity regimes where it surpasses current state-of-the-art methods.
+Code can be found at https://github.com/BaiTheBest/SparseLLM.
 1 Introduction
-Large language models Touvron et al.
+Large language models (LLMs) Touvron et al.
 (2023); OpenAI (2023) have recently transformed the field of natural language processing (NLP) by delivering exceptional results across a variety of intricate language benchmarks Wei et al.
@@
 Nonetheless, these models, with billions of parameters, generally necessitate significant computational resources.
-To make large language models (LLMs) more accessible, extensive efforts have been devoted to model compression of LLMs Xu and McAuley (2023); Ling et al.
-(2023); Bai et al.
-(2024), including pruning, quantization, knowledge distillation, and low-rank factorization, etc.
+To make LLMs more accessible, extensive efforts have been devoted to model compression of LLMs Xu and McAuley (2023); Bai et al.
+(2024), including pruning, quantization, knowledge distillation, and low-rank factorization.
 Pruning, by introducing sparsity, jointly enhances memory and computational efficiency and offers unparalleled flexibility, seamlessly integrating with any LLMs, thus standing out as a highly effective and widely adopted compression strategy.
-Figure 1: AdaGP decomposes the global pruning of LLMs into manageable subproblems by leveraging the chain of modules and auxiliary variables.
 Model pruning has a long history LeCun et al.
@@
 Recently, several local pruning methods, which consider compressing each layer separately, and the overall compressed model is then obtained by “stitching together” the individually compressed layers, are proposed for billion-scale LLMs.
-Frantar and Alistarh (2023) developed SparseGPT, an efficient unstructured pruning method for LLMs with hundreds of billions of parameters, achieving up to 60% parameter reduction with minimal performance loss.
-Sun et al.
-(2023) introduced a novel pruning criterion in Wanda, which evaluates weights by considering both magnitude and related input activations.
-However, despite its efficiency gains, local pruning only aims to minimize the local error for each specific layer under sparsity constraints and hence the final solution is suboptimal to the final goal of the global pruning.
-In other words, beyond the genuine goal of pruning which is just to align the input and output of the pruned and original models, local pruning over-aligns all the intermediate layers’ activations between them, resulting in a suboptimal solution, especially in high-sparsity regimes Singh and Alistarh (2020); Sung et al.
+SparseGPT Frantar and Alistarh (2023), an efficient unstructured pruning method for LLMs with hundreds of billions of parameters, achieves up to 60% parameter reduction with minimal performance loss.
+Another approach, Wanda Sun et al.
+(2023), introduces a novel pruning criterion that evaluates weights by considering both magnitude and related input activations.
+Despite its efficiency gains, local pruning only aims to minimize the local error for each specific layer under sparsity constraints, resulting in a suboptimal solultion for the overall model.
+This is because local pruning over-aligns the intermediate layers’ activations, leading to suboptimal performance, especially in high-sparsity regimes Singh and Alistarh (2020); Sung et al.
 (2023).
-To address these challenges, namely, to achieve global pruning with low memory consumption, we propose Adaptive Global Pruning (AdaGP), that decomposes the global pruning objective into many subproblems, each of which can be solved using low resources and can coordinate each other toward the global pruning objective.
-More specifically, we first formulate LLMs as a composite function where the output of one module is the input of the next, based on which we reformulate the global pruning goal into its equivalent form, with auxiliary variables that facilitate its decomposition and its subproblems’ coordination.
-Then we propose an algorithm that achieves the alternating optimization of the subproblems of all the variables and auxiliary variables with low resource and global optimality, thanks to the closed-form solution of each subproblem.
-Empirically, we find that AdaGP can consistently improve the performance of local pruning methods, particularly in high sparsity regimes (> 60%), where the perplexity can be significantly decreased by up to around 80% as compared to the state-of-the-art local pruning methods, SparseGPT.
-Furthermore, our AdaGP framework can be readily applicable to enhance the performance of most existing local pruning solvers, such as SparseGPT and Wanda, with marginal additional computational overhead.
+Figure 1: SparseLLM decomposes the global pruning of LLMs into manageable subproblems by leveraging the chain of modules and auxiliary variables while maintaining dependencies.
+To address these challenges and achieve global pruning with low memory consumption, we propose SparseLLM that decomposes the global pruning objective into multiple subproblems, each of which can be solved with low resources and coordinate to achieve the global pruning objective.
+More specifically, we first formulate LLMs as a composite function where the output of one module is the input of the next.
+Based on this formulation, we reformulate the global pruning goal into an equivalent form with auxiliary variables that facilitate its decomposition and coordination of the subproblems.
+Then we propose an alternating optimization algorithm to efficiently solve the subproblems, achieving computational resource efficiency and global optimality, due to the close-form solution of each subproblem.
+Empirically, we find that SparseLLM can consistently improve the performance of local pruning methods, particularly in high sparsity regimes (> 60%), where the perplexity can be significantly decreased by up to around 80% as compared to the state-of-the-art methods.
+Furthermore, our SparseLLM framework can be readily applicable to enhance the performance of most existing local pruning solvers, such as SparseGPT and Wanda, with marginal additional computational overhead.
 This adaptability ensures that our framework can be seamlessly integrated into a wide range of LLMs and pruning methods, making it a versatile tool and useful baseline for future research exploiting the sparsity of LLMs.
@@
 This approach typically involves setting a threshold to nullify parameters below it, leading to a model with a non-uniform sparse structure.
-Frantar and Alistarh (2023) developed SparseGPT, an efficient unstructured pruning method for LLMs with hundreds of billions of parameters, achieving up to 60% parameter reduction with minimal performance loss.
-Sun et al.
-(2023) introduced a novel pruning criterion in Wanda, which evaluates weights by considering both magnitude and related input activations.
+SparseGPT Frantar and Alistarh (2023), an efficient unstructured pruning method for LLMs with hundreds of billions of parameters, achieves up to 60% parameter reduction with minimal performance loss.
+A novel pruning criterion is introduced in Wanda Sun et al.
+(2023), which evaluates weights by considering both magnitude and related input activations.
 This approach is beneficial in linear layers of LLMs, helping to identify and remove less significant weights.
-Tuli and Jha (2023) proposes DynaTran, a dynamic inference scheme for pruning activations at runtime, supported by a specially designed ASIC architecture, AccelTran, to enhance transformer inference throughput.
-On the other hand, structured pruning involves the selective removal of groups of weights, where ‘group’ might mean blocks of weights, filters, attention heads, or other structures conducive to hardware acceleration.
+Tuli and Jha Tuli and Jha (2023) proposed DynaTran, a dynamic inference scheme for pruning activations at runtime, supported by a specially designed ASIC architecture, AccelTran, to enhance transformer inference throughput.
+On the other hand, structured pruning involves the selective removal of groups of weights, where “group” might mean blocks of weights, filters, attention heads, or other structures conducive to hardware acceleration.
+Ma et al.
 Ma et al.
@@
 Li et al.
+Li et al.
 (2023) proposed LoSparse, a novel approach combining low-rank and sparse matrix approximations to balance pruning and expressive power.
 Tao et al.
+Tao et al.
 (2023) extended this concept to pruning hidden dimensions in LLMs, including embedding layers and attention heads.
-Kurtic et al.
-(2023) presented ZipLM, a structured compression method for LLMs, optimizing for compression and accuracy while considering specific hardware constraints.
-Our work falls in the category of unstructured pruning of LLMs, where existing methods such as SparseGPT and Wanda only consider an entirely local pruning algorithm and suffer from sub-optimal performance.
+ZipLM Kurtic et al.
+(2023), a structured compression method for LLMs, is proposed to optimize for compression and accuracy while considering specific hardware constraints.
+Our work falls in the category of unstructured pruning of LLMs, where existing methods such as SparseGPT and Wanda only consider an entirely local pruning algorithm and suffer from suboptimal performance.
 We discuss the limitations and challenges of entirely local pruning in Sec.
 3.
-3 Background
+3 Background and notation
 3.1 Global pruning
-Given a pre-trained neural network $f$ and inputs $\mathbf{X}$, global pruning aims to find a global sparsity mask $\mathbf{M}$ and possibly updated weights $\widehat{\mathbf{W}}$ to minimize the global loss between the final outputs of the uncompressed and compressed model:
+Given a pre-trained neural network $f$ with parameter $\mathbf{W}$ and inputs $\mathbf{X}$, global pruning aims to find a global sparsity mask $\mathbf{M}$ and possibly updated weights $\widehat{\mathbf{W}}$ to minimize the global loss $\mathcal{L}$ between the final outputs of the uncompressed and compressed model:
 ${\min}_{\mathbf{M},\widehat{\mathbf{{W}}}}\;\;\mathcal{L}\big{(}f(\mathbf{X};\mathbf{M}\odot\widehat{\mathbf{W}}),f(\mathbf{X};\mathbf{W})\big{)},$ (1)
@@
 In addition to NP-hardness Blumensath and Davies (2008), however, a critical challenge in solving Eq.
-1 for modern billion-scale LLMs is the huge memory requirement which prohibits storing the entire model on a single GPU.
-Although there exists toolkits, e.g., Paszke et al.
-(2019); Smith et al.
-(2022), that enable parallelism across multiple GPUs, such a setup may require resources that are not easily accessible, raising interests in developing other pruning methods that are functional regardless of the availability of computational resources.
+1 is the huge memory cost, as one needs to store the entire model in a single GPU, rendering this method impractical for modern billion-scale LLMs.
 3.2 Local pruning
-Local pruning circumvents the memory issue mentioned above by dividing the full model compression into sub-problems for each layer, $\ell$, and constructing a local loss to measure the $\ell_{2}$-error between the outputs of the uncompressed and compressed layers.
+Local pruning circumvents the memory issue mentioned above by dividing the full model compression into subproblems for each layer and constructing a local loss to measure the $\ell_{2}$-error between the outputs of the uncompressed and compressed layers.
 Hence, the local pruning can be formulated by
@@
 (2022).
-Concretely, this means first choosing a pruning mask $\mathbf{M}$ according to some salient criterion, like the weight magnitude Zhu and Gupta (2017), and then optimizing the remaining un-pruned weights while keeping the mask unchanged.
+Concretely, this means first choosing a pruning mask $\mathbf{M}$ according to some salient criterion, like the weight magnitude Zhu and Gupta (2017), and then optimizing the remaining unpruned weights while keeping the mask unchanged.
 Importantly, once the mask is fixed, Eq.
 2 turns into a linear regression problem that can be easily optimized.
-Figure 2: Illustration of our adaptive-global pruning method compared to conventional global pruning and local pruning.
-We consider a two-layer neural network as an abstraction for simplicity.
-Global pruning (left) is memory prohibitive due to poor scalability.
-Local pruning (mid) considers pruning each layer independently, while inevitably sacrificing performance due to the ignorance of global supervision.
-Our adaptive global pruning (right) achieves global pruning with low memory cost by leveraging auxiliary variables and soft constraints.
 Existing solvers.
@@
 Follow-up works demonstrate that pruning accuracy can be further improved by removing the strict separation between mask selection and weight reconstruction.
-In addition, as we mentioned in Section 2, the work Frantar and Alistarh (2023); Sun et al.
-(2023) further advance in the scalability and pruning criterion of the pruning process.
-3.3 What’s Wrong with Local Pruning?
+More recently, Frantar and Alistarh (2023) developed SparseGPT, an efficient unstructured pruning method for LLMs with hundreds of billions of parameters, achieving up to 60% parameter reduction with minimal performance loss.
+Sun et al.
+(2023) introduced a novel pruning criterion in Wanda, which evaluates weights by considering both magnitude and related input activations.
+3.3 What is wrong with local pruning?
 As shown in Eq.
-2, local pruning focuses on minimizing the error for each specific layer $\ell$ subject to sparsity constraints which are embedded in the mask $\mathbf{W}_{\ell}$ and weights $\widehat{\mathbf{W}}_{\ell}$.
-While the primary goal of the pruning process is to ensure that the input and output of the pruned model align closely with those of the original models, local pruning overly constrains the alignment of input and output of all the intermediate layers between the two models, leading to a performance degradation.
-This results in a solution that is suboptimal with respect to the global pruning problem.
-4 Adaptive Global Pruning for LLMs
-We present our proposed method AdaGP that can address the drawbacks of existing pruning methods by achieving a global pruning with low memory consumption.
-AdaGP decomposes the global pruning into many subproblems, each of which can be solved using low resources and can coordinate each other toward the global pruning objective.
+2, local pruning focuses on minimizing the error for each specific layer $\ell$ subject to sparsity constraints.
+This results in a suboptimal solution with respect to the global pruning problem.
+While the primary goal of pruning is to ensure that the input and output of the pruned model align closely with those of the original models, the local pruning overly constrains the activations of all the intermediate layers between the two models, leading to performance degradation.
+4 SparseLLM: Towards global pruning for LLMs
+We present our proposed method SparseLLM that can address the drawbacks of existing pruning methods by achieving a global pruning with low memory consumption.
+SparseLLM decomposes the global pruning objective into many subproblems, each of which can be solved using low resources and can coordinate each other toward the global pruning objective.
+An overview of SparseLLM on the OPT and LlaMA configurations are shown in Figure 2.
 4.1 Motivation
-The development of AdaGP is motivated by the observation: LLMs can be formulated as a composite function such that the output of one module is the input of the next.
-This allows us to reformulate the global pruning goal into its equivalent form with auxiliary variables that enable the decomposition of the original problem into multiple subproblems, as detailed in Figure 2 and Section 4.2.
-Then we develop a resource-efficient algorithm that achieves global optimality via alternating optimization for each of the subproblems in which we obtain closed-form solutions for each of the steps.
-This has been illustrated in Figure 2 and Section 4.3.
+The development of SparseLLM is motivated by the observation: LLMs can be formulated as a composite function such that the output of one module is the input of the next.
+This allows us to reformulate the global pruning goal into its equivalent form with auxiliary variables that enable the decomposition into multiple subproblems, as detailed in Sec.
+4.2.
+Then we develop a resource-efficient algorithm that achieves the alternating optimization of the subproblems with global optimality, thanks to the close-form solution of each subproblem, as illustrated in Sec.
+4.3.
 4.2 A unified formulation of pruning
 In this section, we present the reformulation of the global pruning problem into an equivalent one by introducing auxiliary variables.
-This reformulation enables us to swiftly shift between two extremes of global pruning and local pruning and the decomposition of the problem into many independent subproblems.
-The key idea behind our reformulation is to decouple the densely parametric parts (linear layers) from the non-parametric parts (activation function, self-attention, and layer norm, etc) using a splitting technique.
-Formally, given an LLM formulated as a composite function with dense layers $\ell\in[L]:=\{1,2,\ldots,L\}$, where $L$ is the total number of dense layers and define $\Omega\subseteq[L-1]$ to be a subset of layers for which we control the extent of globality.
-Note that there can be non-parameteric and potentially nonlinear layers between layers $\ell$ and $\ell+1$ in ${L}$ and we denote those layers by function $\phi_{\ell}(\cdot)$.
-In the reformulation, rather than feeding the output of a dense layer $\ell\in[L]$ directly into a non-parametric and potentially non-linear layer after it, we store the output of layer $\ell$ in a new variable $\mathbf{z}_{\ell}=\mathbf{W}_{\ell}\boldsymbol{a}_{\ell-1}$  [^1] For the sake of simplicity and clearer presentation, the bias term is omitted in the following equations where its exclusion does not lead to confusion., where $\boldsymbol{a}_{\ell-1}$ is the input to layer $\ell$ and can be represented by $\boldsymbol{a}_{\ell-1}=\phi_{\ell-1}(\mathbf{z}_{\ell-1})$.
-Note that by definition, $\boldsymbol{a}_{0}$ is the input data.
-Consequently, we can pose the pruning problem as:
+This reformulation provides a more flexible form and enables the decomposition of the problem into many manageable subproblems.
+The key idea behind our formulation is to decouple the densely parametric parts (linear layers) from non-parametric parts (activation function, self-attention, layer norm, etc) using a splitting technique.
+Rather than feeding the output of the dense linear layer $\mathbf{W}_{\ell}$ directly into the non-parametric and potentially nonlinear layer $\phi_{\ell}$, we store the output of layer $\ell$ in a new variable $\mathbf{z}_{\ell}=\mathbf{W}_{\ell}\boldsymbol{a}_{\ell-1}$  [^1] For the sake of simplicity and clearer presentation, the bias term is omitted in the following equations where its exclusion does not lead to confusion..
+We also represent the output of the non-parametric layer as a vector of activations $\boldsymbol{a}_{\ell}=\phi_{\ell}(\mathbf{z}_{\ell})$.
+We then solve the following problem:
 $\begin{split}&\min_{\{\widehat{\mathbf{{W}}}_{\ell}\},\{\mathbf{M}_{\ell}\},\{\boldsymbol{a}_{\ell}\},\{\mathbf{z}_{\ell}\}}\mathcal{L}(\mathbf{z}_{L},\mathbf{y}),\\ &\text{s.t.
 \;\;}\mathbf{z}_{\ell}=(\mathbf{M}_{\ell}\odot\widehat{\mathbf{W}}_{\ell})\boldsymbol{a}_{\ell-1},\;\forall\;\ell\in[L],\\ &\quad\quad\boldsymbol{a}_{\ell}=\phi_{\ell}(\mathbf{z}_{\ell}),\;\forall\;\ell\in\Omega,\\ &\quad\quad\boldsymbol{a}_{\ell},\mathbf{z}_{\ell}=\boldsymbol{a}_{\ell}^{pre},\mathbf{z}_{\ell}^{pre},\;\forall\;\ell\in[L-1]\backslash\Omega,\end{split}$ (3)
-where $[L-1]\backslash\Omega$ denotes the complement set of $\Omega$ and $\boldsymbol{a}_{\ell}^{pre}$ and $\mathbf{z}_{\ell}^{pre}$ denote the corresponding values of $\boldsymbol{a}_{\ell}$ and $\mathbf{z}_{\ell}$ obtained by running a forward-propagation from the pre-trained model respectively.
-In our proposed reformulation above, the key to understanding its unified nature lies in the interpretation of the set $\Omega$. The versatility of our approach is illustrated in the following remark:
-Remark 4.1.
-For the given LLM and definition of $\Omega$, our formulation can seamlessly treat both global and local pruning as special cases depending on the choices of $\Omega$. Specifically:
-  • When $\Omega=[L-1]$, solving our pruning formulation is equivalent to global pruning, accounting for inter-layer dependencies across the entire network.
-  • When $\Omega=\emptyset$, the formulation simplifies to local pruning, considering each layer independently and ignoring inter-layer dependencies.
-The ability to shift between these two extremes, and potentially any intermediate configuration, demonstrates the flexibility and comprehensiveness of our reformulation.
+where $L$ represents the total number of dense (linear) layers and $[L]=\{1,2,\cdots,L\}$.
+$[L-1]\backslash\Omega$ denotes the complement set of $\Omega$.
+We use $\boldsymbol{a}_{\ell}^{pre}$, $\mathbf{z}_{\ell}^{pre}$ to denote the corresponding intermediate variables’ values of the original dense (i.e., without pruning) pre-trained model.
+$y$ denotes the ground-truth final output of the dense pre-trained model.
+In our proposed formulation above, its unified nature lies in the interpretation and application of the set $\Omega$, which denotes the indices of layers subject to the pruning process.
+Intuitively, $\Omega$ measures how “global” the pruning is.
+The bigger the set of $\Omega$ is, the more layers are connected via the second constraint, and the pruning is more towards the global extreme, and vice versa.
+The generality and versatility of our formulation is illustrated in the following remark:
+Remark 4.1 (Generality and flexibility of Eq. 3).
+Given an LLM formulated as a composite function with dense layers $l\in\{1,2,\ldots,L-1\}$, where $L$ is the total number of dense layers and $\Omega$ denotes the set of layers subject to the pruning process.
+Our formulation can seamlessly treat both global and local pruning as special cases under certain conditions.
+Specifically:
+  • When $\Omega=\{1,2,\ldots,L-1\}$, solving our pruning formulation is equivalent to global pruning, accounting for inter-layer dependencies across the entire network.
+  • When $\Omega=\emptyset$, the formulation simplifies to local pruning, considering each layer independently (the last constraint dominates and “cuts” all layer dependencies with pre-trained values.)
+The ability to shift between these two extremes, and potentially any intermediate configurations, demonstrates the flexibility and comprehensiveness of our formulation.
+By adjusting $\Omega$, one can seamlessly transition from a global perspective to a local perspective.
 This flexibility not only caters to a wide range of pruning strategies but also provides a unified framework to compare and contrast the effectiveness of different pruning methods under a consistent mathematical lens.
+Figure 2: Illustration of SparseLLM on OPT and LlaMA.
+The auxiliary variables and soft constraints (i.e., $\approx$) allow SparseLLM to decompose the global pruning into manageable subproblems while maintaining the dependencies.
+Subproblems are analytically solvable and enjoy fast convergence.
 4.3 Algorithm design
-In this section, we introduce a choice of $\Omega$ that achieves a good trade-off between global and local pruning and propose an algorithm to solve the resulting problem.
-Specifically, we propose an adaptive approach where global pruning is applied within each decoder layer, while still adhering to a “local pruning” strategy between different decoder layers.
-Moreover, given that the feed-forward network (FFN) module in each decoder layer accounts for more than two-thirds of the total parameters Liu et al.
-(2024), our approach prioritizes the global pruning of these FFN layers.
-This strategy strikes a balance between the computational feasibility of pruning large-scale models and the effectiveness of the pruning process.
-Formally, we select $\Omega$ to be the FFN layers, denoted by $\Omega_{\text{FFN}}$.
-As a result, we obtain the problem of interest by substituting $\Omega$ with $\Omega_{\text{FFN}}$ in Eq.
-3.
-We have the following problem now:
-$\begin{split}&\min_{\{\widehat{\mathbf{{W}}}_{\ell}\},\{\mathbf{M}_{\ell}\},\{\boldsymbol{a}_{\ell}\},\{\mathbf{z}_{\ell}\}}\mathcal{L}(\mathbf{z}_{L},\mathbf{y}),\\ &\text{s.t.
-\;}\mathbf{z}_{\ell}=(\mathbf{M}_{\ell}\odot\widehat{\mathbf{W}}_{\ell})\boldsymbol{a}_{\ell-1},\;\forall\;\ell\in[L],\\ &\quad\;\;\,\boldsymbol{a}_{\ell}=\phi_{\ell}(\mathbf{z}_{\ell}),\;\forall\;\ell\in\Omega_{\text{FFN}},\\ &\quad\;\boldsymbol{a}_{\ell},\mathbf{z}_{\ell}=\boldsymbol{a}_{\ell}^{pre},\mathbf{z}_{\ell}^{pre},\;\forall\;\ell\in[L-1]\backslash\Omega_{\text{FFN}},\end{split}$ (4)
-Instead of solving Eq.
-4 directly, we relax the constraints by adding an $\ell_{2}$-penalty function to the objective to obtain an unconstrained problem.
-In addition, by considering the “local pruning" between the decoder layer, we can simplify Eq.
-4 to obtain a unconstrained subproblem for each of the decoder layer as:
-$\begin{split}&\min_{\{\widehat{\mathbf{{W}}}_{\ell}\},\{\mathbf{M}_{\ell}\},\boldsymbol{a}_{\ell},\mathbf{z}_{\ell}}\alpha\|\mathbf{z}_{\ell+1}^{pre}-(\mathbf{M}_{\ell+1}\odot\widehat{\mathbf{W}}_{\ell+1})\boldsymbol{a}_{\ell}\|_{2}^{2}\\ &\quad\quad\quad+\beta\|\boldsymbol{a}_{\ell}-\phi_{\ell}(\mathbf{z}_{\ell})\|_{2}^{2},\\ &\quad\quad\quad+\alpha\|\mathbf{z}_{\ell}-(\mathbf{M}_{\ell}\odot\widehat{\mathbf{W}}_{\ell})\boldsymbol{a}_{\ell-1}^{pre}\|_{2}^{2},\end{split}$ (5)
-where $\alpha$, $\beta$ are hyperparameters for the penalty of corresponding constraint in Eq.
-4 and layers $\ell$ and $\ell+1$ correspond to the up-scaling and down-scaling linear layers in FFN.
-Note that for the last decoder layer, we replace $\mathbf{z}_{\ell+1}^{pre}$ with $y$.
+In this section, we introduce the algorithm design of SparseLLM, which alternatively optimizes the subproblems associated with the corresponding variables.
+This approach is resource-efficient and achieves global optimality, attributed to the closed-form solutions that each subproblem yields.
+The key idea of our algorithm lies behind the flexibility of $\Omega$ in our Eq.
+3, as we want to find a better trade-off between completely global (memory bottleneck) and completely local (suboptimal performance) pruning.
+Naively applying SparseLLM to prune all layers globally is impractical.
+On the other hand, recent work shows that the feed-forward network (FFN) module in each decoder layer accounts for more than two-thirds of the total parameters in an LLM Liu et al.
+(2024).
+Therefore, our SparseLLM prioritizes the global pruning of the FFN module, while still adhering to a local pruning strategy for the multi-head attention (MHA) module (see Figure 2).
+This strategy strikes a balance between the computational feasibility of pruning large-scale models and the effectiveness of the pruning process, adhering to the limitations and practices of state-of-the-art LLM pruning frameworks.
+Formally speaking, rather than try to solve Eq. 3 directly, we first relax the constraints by adding an $\ell_{2}$-penalty function to the objective and attack the unconstrained problem:
+$\begin{split}\mathcal{L}(\mathbf{z}_{L},\mathbf{y})+\alpha\sum\nolimits_{\ell\in[L]}\|\mathbf{z}_{\ell}-(\mathbf{M}_{\ell}\odot\widehat{\mathbf{W}}_{\ell})\boldsymbol{a}_{\ell-1}\|_{2}^{2}+\beta\sum\nolimits_{\ell\in\Omega_{\text{FFN}}}\|\boldsymbol{a}_{\ell}-\phi_{\ell}(\mathbf{z}_{\ell})\|_{2}^{2},\end{split}$ (4)
+where $\alpha$, $\beta$ are hyperparameters for controlling the weight of each constraint.
+$\Omega_{\text{FFN}}$ denotes the set of indexes for the linear layers in the FFN module of each decoder layer, i.e., linear layers from the same FFN module are pruned globally.
+For simplicity, the superscript “pre” of $\boldsymbol{a}_{\ell}$ and $\mathbf{z}_{\ell}$ in the third constraint in Eq.
+3 is omitted here, i.e., for $\ell\not\in\Omega_{\text{FFN}}$ the $\boldsymbol{a}_{\ell}$ and $\mathbf{z}_{\ell}$ are fixed and equal to the pre-trained model’s intermediate value in the second term of Eq.
+4.
+In the following subsections, we illustrate how we approach the pruning of FFN and MHA modules, respectively.
+4.3.1 SparseLLM on OPT models
+For each decoder layer in a pre-trained LLM, our Eq. 4 instantly simplifies to globally pruning the corresponding FFN module within that decoder layer as:
+$\begin{split}\alpha\|\mathbf{z}_{\ell+1}^{pre}-(\mathbf{M}_{\ell+1}\odot\widehat{\mathbf{W}}_{\ell+1})\boldsymbol{a}_{\ell}\|_{2}^{2}+\beta\|\boldsymbol{a}_{\ell}-\phi_{\ell}(\mathbf{z}_{\ell})\|_{2}^{2}+\alpha\|\mathbf{z}_{\ell}-(\mathbf{M}_{\ell}\odot\widehat{\mathbf{W}}_{\ell})\boldsymbol{a}_{\ell-1}^{pre}\|_{2}^{2},\end{split}$ (5)
+where layers $\ell$ and $\ell+1$ correspond to the up-projection and down-projection linear layers.
 In this work, we consider the alternating method to optimize our Eq.
@@
 The careful and elaborate design of our Eq.
-5 allows us to derive a closed-form solution for each of the variables:
+5 allows us to derive a closed-form solution to every subproblem as shown in the following.
 Pruning weight.
@@
 For each linear layer $\ell$ in a FFN module, the optimal solution minimizes $\|\mathbf{z}_{\ell}-(\mathbf{M}_{\ell}\odot\widehat{\mathbf{W}}_{\ell})\boldsymbol{a}_{\ell-1}\|_{2}^{2}$.
-To solve it, the first step is to write $\mathbf{z}_{\ell}$ as $\mathbf{W}_{\ell}\boldsymbol{a}_{\ell-1}$, where $\mathbf{W}_{\ell}=\mathbf{z}_{\ell}\boldsymbol{a}_{\ell-1}^{\dagger}$ ($\dagger$ denotes the pseudo-inverse.).
-We get $\|\mathbf{W}_{\ell}\boldsymbol{a}_{\ell-1}-(\mathbf{M}_{\ell}\odot\widehat{\mathbf{W}}_{\ell})\boldsymbol{a}_{\ell-1}\|_{2}^{2}$ by plugging $\mathbf{z}_{\ell}$ back in original loss.
-This aligns with the pruning objective of Eq.
+To solve it, the first step is to decompose $\mathbf{z}_{\ell}$ to $\mathbf{W}_{\ell}\boldsymbol{a}_{\ell-1}$, where $\mathbf{W}_{\ell}=\mathbf{z}_{\ell}\boldsymbol{a}_{\ell-1}^{\dagger}$ ($\dagger$ denotes the pseudo-inverse.) Plug decomposed $\mathbf{z}_{\ell}$ back in original loss and we get $\|\mathbf{W}_{\ell}\boldsymbol{a}_{\ell-1}-(\mathbf{M}_{\ell}\odot\widehat{\mathbf{W}}_{\ell})\boldsymbol{a}_{\ell-1}\|_{2}^{2}$, which aligns with the pruning objective of Eq.
 2 and can be analytically solved by existing pruning solver e.g., SparseGPT.
@@
 By following a very similar idea to Ridge regression, the new value of $\boldsymbol{a}_{\ell}$ is given by:
-$(\alpha\mathbf{W}_{\ell+1}^{\intercal}\mathbf{W}_{\ell+1}+\beta\mathbf{I})^{-1}(\alpha\mathbf{W}_{\ell+1}^{\intercal}\mathbf{z}_{\ell+1}^{pre}+\beta\phi_{\ell}(\mathbf{z}_{\ell})),$ (6)
+$(\alpha\mathbf{W}_{\ell+1}^{\intercal}\mathbf{W}_{\ell+1}+\beta\mathbf{I})^{-1}(\alpha\mathbf{W}_{\ell+1}^{\intercal}\mathbf{z}_{\ell+1}^{pre}+\beta\cdot\text{ReLU}(\mathbf{z}_{\ell})),$ (6)
 where $\mathbf{W}_{\ell}$ denotes the updated weight matrix after pruning, i.e., $\mathbf{W}_{\ell}\coloneqq\mathbf{M}_{\ell}\odot\widehat{\mathbf{W}}_{\ell}$.
-Updating Output.
-The update for $\mathbf{z}_{\ell}$ requires minimizing the following loss:
-$\beta\|\boldsymbol{a}_{\ell}-\phi_{\ell}(\mathbf{z}_{\ell})\|_{2}^{2}+\alpha\|\mathbf{z}_{\ell}-(\mathbf{M}_{\ell}\odot\widehat{\mathbf{W}}_{\ell})\boldsymbol{a}_{\ell-1}^{pre}\|_{2}^{2}.$ (7)
-This problem can be challenging because of the potentially non-linear function $\phi_{\ell}(\cdot)$.
-Fortunately, as the non-linearity $\phi_{\ell}(\cdot)$ between two consecutive FFN layers in the transformer is typically ReLU, which works entry-wise on its argument, the entries in $\mathbf{z}_{\ell}$ are decoupled.
-As a result, solving Eq.
-7 is particularly easy when $\phi_{\ell}$ is piece-wise linear (as in the case of ReLU).
-It can be solved in closed-form following a simple if-then logic.
-Specifically, one only needs to compute two solutions as:
-$\begin{split}&\mathbf{z}_{\ell}^{(1)}=(\mathbf{M}_{\ell}\odot\widehat{\mathbf{W}}_{\ell})\boldsymbol{a}_{\ell-1}^{pre},\\ &\mathbf{z}_{\ell}^{(2)}=(\alpha+\beta)^{-1}\cdot\big{(}\beta\boldsymbol{a}_{\ell}+\alpha\mathbf{z}_{\ell}^{(1)}\big{)},\end{split}$ (8)
-where the first solution corresponds to those entries of $\mathbf{z}_{\ell}$ that are negative (reduced to zero by ReLU), and the second solution corresponds to those entries of $\mathbf{z}_{\ell}$ that are non-negative (i.e., $\phi_{\ell}(\mathbf{z}_{\ell})=\mathbf{z}_{\ell}$).
-For more sophisticated choices of $\phi_{\ell}$, including smooth sigmoid curves, the problem can be solved quickly with a lookup table of pre-computed solutions because each entry only depends on two inputs.
-Remark 4.2 (Convergence to global optimum with AdaGP).
-Consider the unconstrained problem given in Eq.
-5, we observe that, if the activation function $\phi$ is ReLU, (1) the objective function is convex with respect to each variable when all others are fixed, and (2) closed-form solutions exist for each of the variable updates in the alternating optimization scheme.
-The proposed algorithm resembles a multi-block alternating direction method of multipliers (ADMM) that is shown to converge to the global optimum in many applications.
-4.4 Time Complexity Analysis
-The proposed AdaGP is composed of three main steps, with the overall time complexity being the sum of the complexities of the individual steps.
-Consider one epoch of AdaGP, in the step of weights pruning, the time complexity is dominated by the computation of the pseudo-inverse of matrix $\boldsymbol{a}_{\ell}$ with dimensions $n\times h$ (where $n$ is the number of samples and $h$ is the hidden dimension of the transformer) and thus given by $O(nh^{2})$, assuming $n<h$.
-The exact pruning step, when employing SparseGPT as the solver, has a complexity of $O(h^{3})$.
-Combining both sub-steps, weight pruning has a total time complexity of $O(h^{3})$.
-The second step, updating activation, is dominated by a matrix inversion as shown in Eq.
-6.
-Given that the weight matrix $\mathbf{W}_{\ell}$ is of size $h\times h$, this inversion has a time complexity of $O(h^{3})$.
-The third step, updating outputs, does not involve matrix inversion and its complexity does not exceed that of the previous steps.
-Thus, the overall time complexity of the algorithm is bounded by $O(h^{3})$.
-Consequently, our method exhibits a per-epoch time complexity that is on par with SparseGPT.
+Updating output. The update for $\mathbf{z}_{\ell}$ requires minimizing the following loss:
+$\beta\|\boldsymbol{a}_{\ell}-\text{ReLU}(\mathbf{z}_{\ell})\|_{2}^{2}+\alpha\|\mathbf{z}_{\ell}-(\mathbf{M}_{\ell}\odot\widehat{\mathbf{W}}_{\ell})\boldsymbol{a}_{\ell-1}^{pre}\|_{2}^{2}.$ (7)
+This problem is non-convex and non-quadratic (because of the non-linear function ReLU).
+Fortunately, because the ReLU function works entry-wise on its argument, the entries in $\mathbf{z}_{\ell}$ are de-coupled.
+Solving Eq.
+7 is particularly easy for the case of ReLU, as it can be solved in closed form followed by a simple if-then logic.
+Specifically, one only needs to compute two solutions of a quadratic equation:
+$\begin{split}\mathbf{z}_{\ell}^{(1)}=(\mathbf{M}_{\ell}\odot\widehat{\mathbf{W}}_{\ell})\boldsymbol{a}_{\ell-1}^{pre},\quad\mathbf{z}_{\ell}^{(2)}=(\alpha+\beta)^{-1}\cdot\big{(}\beta\boldsymbol{a}_{\ell}+\alpha\mathbf{z}_{\ell}^{(1)}\big{)},\end{split}$ (8)
+where the first solution corresponds to those entries of $\mathbf{z}_{\ell}$ that are negative (reduced to zero by ReLU), and the second solution corresponds to those entries of $\mathbf{z}_{\ell}$ that are non-negative.
+4.3.2 SparseLLM on LlaMA models
+In this section, we introduce how SparseLLM decomposes global pruning into subproblems and solves them iteratively on LlaMA model families.
+The model architecture of LlaMA can be found in Figure 2.
+Overall, SparseLLM operates similarly on both LlaMA and OPT models, with the main difference being that LlaMA includes an additional dense linear layer, known as the gate projection layer, and uses the SiLU activation function instead of ReLU.
+Pruning weight. In this part, SparseLLM functions almost identically to its operation on OPTs.
+Updating activation $\mathbf{a}_{\ell}$.
+Similarly, for updating $\boldsymbol{a}_{\ell}$, SparseLLM works nearly the same as on OPT.
+The minimization for $\boldsymbol{a}_{\ell}$ is a simple least-squares problem, akin to weight pruning.
+However, in this case, the matrix $\boldsymbol{a}_{\ell-1}$ appears in two penalty terms in Eq.
+5, necessitating the minimization of:
+$\alpha\|\mathbf{z}_{\ell+1}^{pre}-(\mathbf{M}_{\ell+1}\odot\widehat{\mathbf{W}}_{\ell+1})\boldsymbol{a}_{\ell}\|_{2}^{2}+\beta\|\boldsymbol{a}_{\ell}-\text{SiLU}(\mathbf{s}_{\ell})\odot\mathbf{z}_{\ell}\|_{2}^{2},$ (9)
+for $\boldsymbol{a}_{\ell}$, with all other variables held fixed. Following a concept similar to Ridge regression, the updated value of $\boldsymbol{a}_{\ell}$ is:
+$\big{(}\alpha\mathbf{W}_{\ell+1}^{\intercal}\mathbf{W}_{\ell+1}+\beta\mathbf{I}\big{)}^{-1}\big{(}\alpha\mathbf{W}_{\ell+1}^{\intercal}\mathbf{z}_{\ell+1}^{pre}+\beta\cdot\text{SiLU}(\mathbf{s}_{\ell})\odot\mathbf{z}_{\ell}\big{)},$ (10)
+where $\mathbf{W}_{\ell}$ denotes the updated weight matrix after pruning, i.e., $\mathbf{W}_{\ell}\coloneqq\mathbf{M}_{\ell}\odot\widehat{\mathbf{W}}_{\ell}$.
+Updating output $\mathbf{z}_{\ell}$.
+Updating $\mathbf{z}_{\ell}$ is somewhat simpler in LlaMA since the activation function applies over the gate projection layer.
+The update requires minimizing the loss:
+$\beta\|\boldsymbol{a}_{\ell}-\text{SiLU}(\mathbf{s}_{\ell})\odot\mathbf{z}_{\ell}\|_{2}^{2}+\alpha\|\mathbf{z}_{\ell}-(\mathbf{M}_{\ell}\odot\widehat{\mathbf{W}}_{\ell})\boldsymbol{a}_{\ell-1}^{pre}\|_{2}^{2}.$ (11)
+This problem is quadratic when solving for $\mathbf{z}_{\ell}$ with other variables fixed.
+Through mathematical manipulations, the analytical solution for $\mathbf{z}_{\ell}$ is found by solving a quadratic equation:
+$\mathbf{z}_{\ell}^{*}=\frac{(\mathbf{M}_{\ell}\odot\widehat{\mathbf{W}}_{\ell})\boldsymbol{a}_{\ell-1}^{pre}+\text{SiLU}(\mathbf{s}_{\ell})\odot\boldsymbol{a}_{\ell}}{\text{SiLU}(\mathbf{s}_{\ell})^{2}+\mathbf{1}},$ (12)
+where the division is element-wise and $\mathbf{1}$ denotes the all-one matrix.
+Updating gate projection output $\mathbf{s}_{\ell}$. Updating $\mathbf{s}_{\ell}$ involves minimizing:
+$\beta\|\boldsymbol{a}_{\ell}-\text{SiLU}(\mathbf{s}_{\ell})\odot\mathbf{z}_{\ell}\|_{2}^{2}+\alpha\|\mathbf{s}_{\ell}-(\mathbf{M}_{s}\odot\widehat{\mathbf{W}}_{s})\boldsymbol{a}_{\ell-1}^{pre}\|_{2}^{2},$ (13)
+where $\mathbf{M}_{s}$ and $\widehat{\mathbf{W}}_{s}$ denote the mask and layer weights for the gate projection layer.
+This problem is non-convex and non-quadratic due to the non-linear SiLU function.
+However, since SiLU operates entry-wise, the entries in $\mathbf{s}_{\ell}$ are decoupled.
+Despite LlaMA lacking a simple closed-form solution as in OPT (which uses ReLU), the problem can still be solved quickly and analytically using a lookup table of pre-computed solutions, since each element in $\mathbf{s}_{\ell}$ depends on only three variables.
+Remark 4.2 (Global convergence of SparseLLM).
+Consider the objective function given by Eq.
+5, under the condition that the activation function $\phi$ is ReLU.
+Notice that (1) the objective function is convex with respect to each variable when all others are fixed, and (2) given that closed-form solutions exist for the subproblems in the alternating optimization scheme, the proposed algorithm is guaranteed to converge to a global optimum.
+4.3.3 Pruning of MHAs
+SparseLLM also prunes other linear layers besides those in FFNs.
+By following Eq.
+4, for each linear layer out of FFN modules, the pruning objective simplifies to $\alpha\|\mathbf{z}_{\ell+1}^{pre}-(\mathbf{M}_{\ell+1}\odot\widehat{\mathbf{W}}_{\ell+1})\boldsymbol{a}_{\ell}^{pre}\|_{2}^{2}$, which is equivalent (with some simple math) to that of completely local pruning as shown in Eq.
+2.
+Existing LLM pruning solvers such as SparseGPT and Wanda are applicable here.
+4.4 Time complexity analyses
+The proposed SparseLLM consists of three main steps, with the overall time complexity being the sum of the complexities of these steps.
+In the weights pruning step, the complexity is dominated by the pseudo-inverse computation of matrix $\boldsymbol{a}_{\ell}$ (dimensions $n\times h$), which is $O(nh^{2})$.
+Using SparseGPT as the solver, the exact pruning step has a complexity of $O(h^{3})$.
+The second step, updating activation, involves matrix inversion of the weight matrix $\mathbf{W}_{\ell}$ (size $h\times h$) with a complexity of $O(h^{3})$.
+The third step, updating outputs, has a lower complexity.
+Thus, the overall algorithm complexity is bounded by $O(h^{3})$, making our method’s per-epoch time complexity comparable to SparseGPT.
 5 Experiments
 Experiments setup.
-We implement AdaGP in PyTorch Paszke et al.
+We implement SparseLLM in PyTorch Paszke et al.
 (2019) and use the HuggingFace Transformers library Wolf et al.
@@
 All pruning experiments are conducted on NVIDIA A100 GPUs.
-Similar to Frantar and Alistarh (2023); Sun et al.
-(2023), we sparsify each transformer decoder layer sequentially in order, which significantly reduces memory requirements, and only apply global pruning to FFN modules.
-For the calibration data, we follow Frantar and Alistarh (2023) and use 2048-token segments, randomly chosen from the first shard of the C4 Raffel et al.
+For calibration data, we follow Frantar and Alistarh (2023) and use 128 2048-token segments, randomly chosen from the first shard of the C4 Raffel et al.
 (2020) dataset.
 This represents generic text data crawled from the internet and ensures our experiments remain zero-shot as no task-specific data is seen during pruning.
+We followed existing work Frantar and Alistarh (2023); Sun et al.
+(2023) and pruned all linear layers (in FFN and MHA) to the target sparsity.
 Models, datasets & evaluation.
-Following Frantar and Alistarh (2023), we primarily work with the OPT model family Zhang et al.
-(2022) to study scaling behavior at various model sizes, but our method can be easily applied to other LLMs such as Llama Touvron et al.
-(2023).
-While our focus lies on the very largest variants, we also show some results on smaller models to provide a broader picture.
+We consider the OPT model family Zhang et al.
+(2022) and LlaMA-2 model family Touvron et al.
+(2023) in our experiments.
+We show results on different sizes of models to provide a broader picture for the performances of SparseLLM.
 In terms of metrics, we mainly focus on perplexity, which is known to be a challenging and stable metric that is well-suited for evaluating the accuracy of compression methods Yao et al.
@@
 We consider the test sets of raw-WikiText2 Merity et al.
-(2016) and PTB Marcus et al.
+(2016) (WT2) and PTB Marcus et al.
 (1994) as well as a subset of the C4 validation data, all popular benchmarks in LLM compression literature Yao et al.
@@
 (2023).
-OPT-1.3b (WikiText2: 14.62; PTB: 20.29; C4: 16.07)
-Sparsity
-70% 80% 90% 2:4
-Dataset
-WikiText2
-PTB
-C4
-WikiText2
-PTB
-C4
-WikiText2
-PTB
-C4
-WikiText2
-PTB
-C4
-Magnitude
-6420.80
-4828.13
-3435.99
-9998.71
-1.1e4
-5347.89
-8209.13
-1.0e4
-4917.02
-96.68
-133.92
-48.08
-Wanda
-21.56
-34.77
-25.78
-142.20
-146.76
-142.24
-5692.65
-4751.69
-4501.73
-15.63
-24.04
-18.23
-SparseGPT
-18.04
-28.19
-21.45
-69.67
-93.36
-60.83
-2596.70
-2361.86
-1363.08
-15.11
-23.71
-17.88
-AdaGP
-17.82
-27.72
-20.99
-58.92
-85.33
-58.36
-1350.31
-1192.36
-655.76
-14.97
-23.40
-17.67
-OPT-2.7b (WikiText2: 12.47; PTB: 17.97; C4: 14.32)
-Sparsity
-70% 80% 90% 2:4
-Dataset
-WikiText2
-PTB
-C4
-WikiText2
-PTB
-C4
-WikiText2
-PTB
-C4
-WikiText2
-PTB
-C4
-Magnitude
-1691.74
-1237.08
-1415.02
-1.0e4
-7916.69
-6050.07
-7.9e5
-5.3e5
-4.7e5
-272.34
-308.55
-267.70
-Wangda
-88.61
-140.09
-90.06
-6140.81
-4746.96
-5678.66
-3.0e4
-3.5e4
-2.4e4
-13.66
-21.67
-16.10
-SparseGPT
-13.79
-21.18
-16.18
-24.32
-37.82
-25.92
-2662.74
-2285.01
-1776.08
-12.62
-19.28
-15.12
-AdaGP
-13.82
-21.07
-16.14
-23.87
-37.09
-24.90
-1200.12
-759.11
-527.70
-12.63
-19.31
-15.12
-OPT-6.7b (WikiText2: 10.86; PTB: 15.77; C4: 12.71)
-Sparsity
-70% 80% 90% 2:4
-Dataset
-WikiText2
-PTB
-C4
-WikiText2
-PTB
-C4
-WikiText2
-PTB
-C4
-WikiText2
-PTB
-C4
-Magnitude
-7054.21
-5437.44
-4850.25
-7937.49
-5971.86
-6031.54
-2.4e4
-2.5e4
-2.1e4
-64.11
-92.23
-82.67
-Wanda
-54.95
-129.73
-116.67
-1493.58
-1196.93
-996.00
-2.1e4
-2.0e4
-1.8e4
-11.86
-18.54
-14.77
-SparseGPT
-12.27
-18.90
-15.28
-31.04
-51.26
-29.42
-8871.24
-5713.57
-3797.20
-11.30
-16.90
-13.51
-AdaGP
-12.16
-18.39
-14.93
-23.96
-39.32
-26.97
-2095.85
-1842.48
-953.44
-11.07
-16.73
-13.42
-OPT-13b (WikiText2: 10.13; PTB: 14.52; C4: 12.06)
-Sparsity
-70% 80% 85% 2:4
-Dataset
-WikiText2
-PTB
-C4
-WikiText2
-PTB
-C4
-WikiText2
-PTB
-C4
-WikiText2
-PTB
-C4
-Magnitude
-9037.12
-7734.58
-5909.47
-1.1e4
-9140.88
-6340.22
-1.3e4
-1.3e4
-9087.50
-67.07
-110.77
-52.61
-Wanda
-30.94
-39.26
-33.31
-4216.04
-2894.77
-2450.57
-1.1e4
-1.1e4
-7244.96
-10.33
-15.35
-12.54
-SparseGPT
-10.89
-16.35
-13.39
-21.42
-33.62
-21.01
-8408.03
-6380.30
-3416.23
-10.20
-15.14
-12.48
-AdaGP
-10.96
-16.57
-13.38
-19.07
-28.77
-19.29
-2052.27
-1536.51
-538.61
-10.21
-15.15
-12.48
-OPT-30b (WikiText2: 9.56; PTB: 14.04; C4: 11.45)
-Sparsity
-70% 80% 90% 2:4
-Dataset
-WikiText2
-PTB
-C4
-WikiText2
-PTB
-C4
-WikiText2
-PTB
-C4
-WikiText2
-PTB
-C4
-Magnitude
-8691.40
-4769.89
-4732.66
-8941.81
-5292.98
-5092.26
-3.8e7
-3.0e7
-1.4e7
-517.71
-374.54
-991.95
-Wanda
-7766.61
-5547.45
-5741.74
-8770.33
-6020.70
-7132.20
-6354.15
-4296.37
-4654.27
-10.94
-16.64
-13.50
-SparseGPT
-9.58
-14.41
-11.93
-16.49
-22.01
-17.67
-5747.87
-5169.50
-3555.24
-9.59
-14.05
-11.60
-AdaGP
-9.56
-14.40
-11.94
-15.61
-19.64
-16.61
-3050.63
-2712.39
-1758.63
-9.57
-14.04
-11.60
-OPT-66b (WikiText2: 9.34; PTB: 13.36; C4: 10.99)
-Sparsity
-70% 80% 90% 2:4
-Dataset
-WikiText2
-PTB
-C4
-WikiText2
-PTB
-C4
-WikiText2
-PTB
-C4
-WikiText2
-PTB
-C4
-Magnitude
-OOM
-OOM
-OOM
-OOM
-OOM
-OOM
-OOM
-OOM
-OOM
-OOM
-OOM
-OOM
-Wanda
-OOM
-OOM
-OOM
-OOM
-OOM
-OOM
-OOM
-OOM
-OOM
-OOM
-OOM
-OOM
-SparseGPT
-9.45
-13.64
-11.37
-28.27
-57.41
-26.26
-7803.10
-6594.88
-4433.35
-9.38
-13.37
-11.12
-AdaGP
-9.37
-13.66
-11.37
-16.45
-21.00
-17.70
-7504.17
-5644.65
-3683.91
-9.36
-13.36
-11.12
-Table 1: Perplexity of pruned OPT models in high sparsity regimes (> 60%); the lower the perplexity, the better.
-OPT-6.7b
-Sparsity
-Method
-BoolQ
-RTE
-HellaSwag
-WinoGrande
-ARC-e
-ARC-c
-OBQA
-Mean
-Dense
-66.12
-56.03
-50.49
-65.27
-65.72
-30.63
-27.60
-51.69
-70%
-SparseGPT
-61.74
-54.87
-48.46
-63.85
-64.31
-29.27
-25.40
-49.70
-AdaGP
-60.61
-54.51
-48.8
-62.9
-64.14
-30.03
-26.60
-49.66
-80%
-SparseGPT
-55.08
-48.38
-42.22
-59.43
-57.79
-25.85
-21.40
-44.31
-AdaGP
-58.69
-51.26
-43.78
-59.67
-58.38
-26.88
-22.00
-45.81
-90%
-SparseGPT
-38.53
-53.07
-26.00
-48.07
-26.81
-21.67
-14.40
-32.65
-AdaGP
-46.48
-52.71
-26.21
-51.70
-27.44
-19.71
-13.40
-33.95
-2:4
-SparseGPT
-62.11
-54.87
-49.58
-64.4
-65.32
-30.63
-27.60
-50.64
-AdaGP
-63.82
-55.96
-49.55
-64.56
-65.70
-30.55
-27.60
-51.11
-OPT-13b
-Sparsity
-Method
-BoolQ
-RTE
-HellaSwag
-WinoGrande
-ARC-e
-ARC-c
-OBQA
-Mean
-Dense
-65.87
-57.76
-52.44
-66.02
-67.82
-33.46
-28.62
-53.14
-70%
-SparseGPT
-63.03
-54.87
-50.89
-65.43
-67.47
-32.85
-26.40
-51.56
-AdaGP
-63.85
-55.23
-50.73
-65.67
-66.46
-31.83
-27.20
-51.57
-80%
-SparseGPT
-59.72
-52.35
-46.82
-61.48
-62.50
-31.23
-21.80
-47.99
-AdaGP
-60.89
-53.07
-46.19
-62.12
-62.21
-30.38
-23.00
-48.27
-90%
-SparseGPT
-47.49
-52.71
-33.17
-51.54
-39.98
-21.33
-17.80
-37.72
-AdaGP
-53.43
-52.71
-38.19
-52.96
-46.68
-25.26
-17.40
-40.95
-2:4
-SparseGPT
-64.71
-54.51
-51.92
-65.98
-67.47
-32.85
-28.20
-52.23
-AdaGP
-65.63
-55.23
-51.93
-65.27
-67.8
-33.45
-28.40
-52.53
+Table 1: Perplexity in high sparsity regimes ($\geq 70\%$); the lower the perplexity, the better.
+OPT-1.3b (WikiText2 (WT2): 14.62; PTB: 20.29; C4: 16.07)
+Sparsity 70% 80% 90% 3:4
+Dataset WT2 PTB C4 WT2 PTB C4 WT2 PTB C4 WT2 PTB C4
+Magnitude 6420.80 4828.13 3435.99 9998.71 1.1e4 5347.89 8209.13 1.0e4 4917.02 - - -
+Wanda 21.56 34.77 25.78 142.20 146.76 142.24 5692.65 4751.69 4501.73 - - -
+SparseGPT 18.04 28.19 21.45 69.67 93.36 60.83 2596.70 2361.86 1363.08 252.81 238.41 146.21
+SparseLLM 17.82 27.72 20.99 58.92 85.33 58.36 1350.31 1192.36 655.76 128.83 144.48 106.01
+OPT-2.7b (WikiText2 (WT2): 12.47; PTB: 17.97; C4: 14.32)
+Sparsity 70% 80% 90% 3:4
+Dataset WT2 PTB C4 WT2 PTB C4 WT2 PTB C4 WT2 PTB C4
+Magnitude 1691.74 1237.08 1415.02 1.0e4 7916.69 6050.07 7.9e5 5.3e5 4.7e5 - - -
+Wanda 88.61 140.09 90.06 6140.81 4746.96 5678.66 3.0e4 3.5e4 2.4e4 - - -
+SparseGPT 13.79 21.18 16.18 24.32 37.82 25.92 2662.74 2285.01 1776.08 91.02 91.79 64.95
+SparseLLM 13.82 21.07 16.14 23.87 37.09 24.90 1200.12 759.11 527.70 56.90 77.14 52.77
+OPT-30b (WikiText2 (WT2): 9.56; PTB: 14.04; C4: 11.45)
+Sparsity 70% 80% 90% 3:4
+Dataset WT2 PTB C4 WT2 PTB C4 WT2 PTB C4 WT2 PTB C4
+Magnitude 8691.40 4769.89 4732.66 8941.81 5292.98 5092.26 3.8e7 3.0e7 1.4e7 - - -
+Wanda 7766.61 5547.45 5741.74 8770.33 6020.70 7132.20 6354.15 4296.37 4654.27 - - -
+SparseGPT 9.58 14.41 11.93 16.49 22.01 17.67 5747.87 5169.50 3555.24 441.35 464.73 209.44
+SparseLLM 9.56 14.40 11.94 15.61 19.64 16.61 3050.63 2712.39 1758.63 51.28 73.61 37.99
+OPT-66b (WikiText2 (WT2): 9.34; PTB: 13.36; C4: 10.99)
+Sparsity 70% 80% 90% 3:4
+Dataset WT2 PTB C4 WT2 PTB C4 WT2 PTB C4 WT2 PTB C4
+Magnitude OOM OOM OOM OOM OOM OOM OOM OOM OOM - - -
+Wanda OOM OOM OOM OOM OOM OOM OOM OOM OOM - - -
+SparseGPT 9.45 13.64 11.37 28.27 57.41 26.26 7803.10 6594.88 4433.35 6594.37 6329.59 3799.87
+SparseLLM 9.37 13.66 11.37 16.45 21.00 17.70 7504.17 5644.65 3683.91 4641.8 5296.93 1618.43
+LlaMA-2 7b (WikiText2 (WT2): 5.47; PTB: 37.91; C4: 7.26)
+Sparsity 70% 80% 90% 3:4
+Dataset WT2 PTB C4 WT2 PTB C4 WT2 PTB C4 WT2 PTB C4
+Magnitude 1058.00 544.43 889.46 6380.27 NaN 4162.92 9498.91 1.02e4 7539.65 - - -
+Wanda 2644.22 4040.95 1630.09 1814.01 3376.35 1124.26 5206.93 4607.30 2780.45 - - -
+SparseGPT 15.98 302.15 18.58 53.20 803.02 52.57 344.97 2503.82 279.77 68.28 784.79 60.45
+SparseLLM 16.15 274.35 18.23 49.96 664.39 47.39 225.23 2233.52 181.56 64.17 667.27 54.56
+LlaMA-2 13b (WikiText2 (WT2): 4.88; PTB: 50.94; C4: 6.73)
+Sparsity 70% 80% 90% 3:4
+Dataset WT2 PTB C4 WT2 PTB C4 WT2 PTB C4 WT2 PTB C4
+Magnitude 30.34 2317.39 28.48 4133.98 4706.65 4112.69 5580.71 5514.22 5090.63 - - -
+Wanda 23.42 502.53 32.65 295.29 2340.13 261.15 3003.49 3804.69 1738.73 - - -
+SparseGPT 12.98 267.63 15.95 45.59 550.59 45.20 825.99 1410.46 673.33 63.48 660.70 56.29
+SparseLLM 12.95 277.76 15.77 36.36 578.35 38.63 646.15 1078.94 466.98 53.71 632.11 50.40
+Table 2: Accuracy (%) of zero-shot tasks; the higher the accuracy, the better.
 OPT-30b
-Sparsity
-Method
-BoolQ
-RTE
-HellaSwag
-WinoGrande
-ARC-e
-ARC-c
-OBQA
-Mean
-Dense
-70.46
-61.82
-54.27
-69.02
-70.47
-35.49
-30.20
-55.96
-70%
-SparseGPT
-68.78
-58.48
-53.83
-67.64
-69.15
-34.30
-29.60
-54.54
-AdaGP
-69.11
-61.73
-53.97
-68.43
-69.78
-34.73
-29.80
-55.36
-80%
-SparseGPT
-64.86
-60.65
-49.73
-61.40
-61.91
-31.74
-24.20
-50.64
-AdaGP
-65.41
-59.57
-50.65
-61.96
-62.71
-32.25
-26.50
-51.29
-90%
-SparseGPT
-37.83
-53.79
-25.96
-49.88
-26.47
-20.22
-12.60
-32.39
-AdaGP
-43.55
-52.35
-26.32
-50.04
-27.31
-20.56
-14.00
-33.45
-2:4
-SparseGPT
-69.54
-58.12
-54.08
-67.25
-69.95
-34.81
-29.80
-54.79
-AdaGP
-69.67
-58.12
-54.08
-68.88
-70.45
-35.49
-30.20
-55.27
-Table 2: Accuracy (%) of zero-shot tasks; the higher the accuracy, the better.
+Sparsity Method BoolQ RTE HellaSwag WinoGrande ARC-e ARC-c OBQA Mean
+Dense 70.46 61.82 54.27 69.02 70.47 35.49 30.20 55.96
+70% SparseGPT 68.78 58.48 53.83 67.64 69.15 34.30 29.60 54.54
+SparseLLM 69.11 61.73 53.97 68.43 69.78 34.73 29.80 55.36
+80% SparseGPT 64.86 60.65 49.73 61.40 61.91 31.74 24.20 50.64
+SparseLLM 65.41 59.57 50.65 61.96 62.71 32.25 26.50 51.29
+90% SparseGPT 37.83 53.79 25.96 49.88 26.47 20.22 12.60 32.39
+SparseLLM 43.55 52.35 26.32 50.04 27.31 20.56 14.00 33.45
+3:4 SparseGPT 55.81 51.26 33.64 54.54 42.05 21.33 21.00 39.95
+SparseLLM 60.83 54.15 39.35 55.41 45.24 24.06 22.20 43.03
+LlaMA-2 7b
+Sparsity Method BoolQ RTE HellaSwag WinoGrande ARC-e ARC-c OBQA Mean
+Dense 75.05 66.43 56.92 69.93 75.34 41.89 34.40 59.99
+70% SparseGPT 68.26 57.04 39.67 59.04 60.9 28.58 20.60 47.73
+SparseLLM 67.61 57.31 40.12 61.39 59.39 28.76 21.40 48.13
+80% SparseGPT 59.36 52.71 28.83 48.7 34.22 18.34 14.40 36.65
+SparseLLM 60.12 53.07 28.62 50.59 34.55 18.69 14.30 37.13
+90% SparseGPT 39.02 52.34 26.66 47.80 28.32 17.37 12.40 31.99
+SparseLLM 39.45 52.71 26.79 51.17 28.32 19.52 12.50 32.92
+3:4 SparseGPT 53.94 54.15 28.09 49.17 31.57 17.41 14.80 35.59
+SparseLLM 57.34 53.43 28.26 48.86 32.45 18.17 14.4 36.13
+LlaMA-2 13b
+Sparsity Method BoolQ RTE HellaSwag WinoGrande ARC-e ARC-c OBQA Mean
+Dense 77.89 70.40 59.94 72.77 77.40 46.50 33.20 62.59
+70% SparseGPT 70.03 53.43 42.20 66.54 64.94 31.66 25.40 50.60
+SparseLLM 69.87 54.15 42.50 68.64 64.97 31.40 25.80 51.05
+80% SparseGPT 62.69 52.71 28.94 50.91 36.24 18.17 14.00 37.67
+SparseLLM 64.39 52.86 29.19 51.46 35.69 18.77 14.20 38.08
+90% SparseGPT 50.21 51.35 26.71 49.14 26.68 19.71 13.2 33.86
+SparseLLM 55.35 52.05 26.89 51.34 27.35 19.62 14.20 35.26
+3:4 SparseGPT 61.28 53.71 28.40 47.99 33.21 18.26 14.00 36.69
+SparseLLM 61.71 55.71 28.56 51.62 32.11 18.49 13.8 37.43
 Comparison methods.
-We compare against the standard magnitude pruning baseline Zhu and Gupta (2017) applied locally.
-We also compare with two state-of-the-art local pruning methods, SparseGPT Frantar and Alistarh (2023) and Wanda Sun et al.
+We compare against three baselines, magnitude pruning Zhu and Gupta (2017) applied locally, and two other state-of-the-art local pruning methods, SparseGPT Frantar and Alistarh (2023) and Wanda Sun et al.
 (2023).
 5.1 Results and analyses
-Pruning vs. Model Sizes.
-We begin by exploring the pruning capabilities of AdaGP across various model sizes in comparison to the baseline methods.
-For each model, we consider unstructured sparsity ranging from 70% to 90% with a 10% increment, as well as a full 2:4 semi-structured sparsity.
+Pruning vs.
+model sizes.
+We begin by exploring the pruning capabilities of SparseLLM across various model sizes in comparison to baseline methods.
+For each model, we consider unstructured sparsity ranging from 70% to 90% with a 10% increment, as well as a 3:4 semi-structured sparsity.
+The 3:4 semi-structured sparsity is inspired by our preliminary results that suggest good performance SparseLLM at high sparsity regimes.
+However, note that two of our baselines, Magnitude and Wanda, are unable to be configured to this sparsity out-of-box.
 We conduct a sensitivity study on the calibration sample sizes (see Appendix A.3) and use calibration sample sizes between 32 and 64 for all experiments.
-Moreover, except for the OPT-125m model, we prune the first 50% of the transformer decoder layers in each model to achieve a balance between the computation resources and showcasing the performances.
+Moreover, we prune the first 50% of the Transformer decoder layers in each model to achieve a balance between the computation resources and the performances.
 Detailed results can be found in Table 1 and Table 5 (see Appendix A.5).
-The perplexity results on the three datasets for the un-pruned dense models are reported next to the names of the models.
-Note that we exclude the results for OPT-350m and adjust 90% unstructured sparsity for OPT-1.3b to 85% unstructured sparsity for numerical stability.
-We provide a discussion on the selection of hyperparameters of $\alpha$ and $\beta$ (used in Eq.
-5) in Appendix A.5.
-From the table, it shows that perplexity increases with increasing sparsity.
-Moreover, we observe a trend of decreasing perplexity for SparseGPT and AdaGP at the same sparsity with increasing model sizes.
+The perplexity results of the dense models are reported next to the names of the models.
+Figure 3: Fast convergence of SparseLLM. Training loss per epoch for pruning layer 3 of OPT-125m at 80% sparsity (Left) and layer 6 of LlaMA-2 13b at 70% sparsity (Right).
+From the table, it shows a general trend of increasing perplexity with increasing sparsity.
+Moreover, we observe a trend of decreasing perplexity for SparseGPT and SparseLLM at the same sparsity with increasing model sizes.
 However, such a trend is not obvious for Magnitude and Wanda.
-We also observe that SparseGPT and AdaGP consistently outperform Magnitude and Wanda by a significant margin.
-For smaller sparsity, AdaGP achieves comparable perplexity to SparseGPT.
-As we increase the sparsity, AdaGP starts to demonstrate noticeable improvements over SparseGPT.
-In numerous instances, AdaGP achieves perplexity reductions of more than 50% compared to SparseGPT.
-Computation Time vs. Model Sizes.
-Next, we study how the computation time for a single epoch of SparseGPT and AdaGP varies with different model sizes, as illustrated in Figure 3.
-We observe that the time taken for a single epoch increases with increasing model sizes.
-The rate at which the time taken increases is comparable for SparseGPT and AdaGP.
-Additionally, although SparseGPT executes only one epoch for each layer to prune, whereas AdaGP may execute multiple epochs depending on the parameter configuration, we find that in most cases, few epochs (less than 10) of AdaGP suffice.
-This suggests that the proposed AdaGP remains computationally efficient.
-Figure 3: Time for one step of SparseGPT and AdaGP.
+We also observe that SparseGPT and SparseLLM consistently outperform Magnitude and Wanda by a significant margin.
+For smaller sparsity, SparseLLM achieves comparable perplexity to SparseGPT.
+As we increase the sparsity, SparseLLM starts to demonstrate noticeable improvements over SparseGPT.
+In numerous instances for the OPT model family, SparseLLM achieves perplexity reductions of more than 50% compared to SparseGPT.
+We also see that performance improvements from SparseLLM over SparseGPT are more significant for the OPT model family than the LlaMA-2 model family.
 Zero-shot experiments.
-To conclude the evaluations and discussions, we show the results for several zero-shot tasks in Table 2 comparing SparseGPT and AdaGP and three model sizes.
+To further conclude the evaluations and discussions, we show results for several zero-shot tasks in Table 2 and Table 6 (see Appendix A.5) comparing SparseGPT and SparseLLM.
 These evaluations are known to be relatively noisy Dettmers et al.
 (2022), but more interpretable.
-We also report the results for zero-shot tasks from the un-pruned dense models in the “Dense" row.
+We also report the results for zero-shot tasks from the dense models in the “Dense" row.
 We see that the accuracy of both methods decreases with increasing sparsity, which is expected, as more parameters are pruned.
 A similar trend of increasing accuracy with increasing model size is observed too.
-Across all the tasks, OBQA and ARC-c remain the most challenging ones as the accuracy for both methods remain around or below 30% while both methods perform well for BoolQ, RTE, WinoGrande, and ARC-e.
-In general, AdaGP is able to achieve higher accuracy in almost all tasks across the three models of different sizes.
+Across all the tasks, OBQA and ARC-c remain the most challenging ones as the accuracy for both methods is 30% or below 30% while both methods perform well for BoolQ, RTE, WinoGrande, and ARC-e.
+In general, SparseLLM is able to achieve higher accuracy in the majority of tasks across the models of different sizes in both OPT and LlaMA-2 model families.
+Training loss vs.
+epochs in SparseLLM.
+Figure 3 illustrates the change in training loss over epochs for SparseLLM, with the training loss plotted on a scale of $10^{3}$ for clarity.
+We observe that the training loss decreases rapidly during the initial epochs, highlighting the efficiency of SparseLLM in achieving effective global pruning within a short period.
+This rapid convergence is largely due to the closed-form solutions employed by SparseLLM for various subproblems, which streamline the pruning process and ensure optimal layer-wise pruning without extensive iterative computations.
+These analytical solutions enable SparseLLM to perform precise pruning operations quickly, making it a powerful tool for optimizing large-scale models like LlaMA, significantly reducing model size while maintaining high accuracy.
 6 Conclusion
-Our work presents AdaGP, a cutting-edge framework poised to redefine the compression of LLMs through sparsity.
-By adeptly circumventing the scalability issues of global pruning and addressing the local suboptimality of existing methods, AdaGP stands as a significant advancement in the field.
+Our work presents SparseLLM, a cutting-edge framework poised to redefine the compression of LLMs through sparsity.
+By adeptly circumventing the scalability issues of global pruning and optimizing the local suboptimality of existing methods, SparseLLM stands as a significant advancement in the field.
 Our empirical results affirm its efficacy, particularly in high-sparsity environments, where it achieves a notable reduction in perplexity, thereby setting a new precedent for model compression.
-The versatility and minimal computational overhead of AdaGP complement its integration with current pruning technologies, underscoring its potential as a universal tool for enhancing the performance and accessibility of LLMs.
-7 Limitations
-While AdaGP marks a significant step forward in efficient pruning of large language models, it is important to acknowledge the inherent trade-off associated with any model compression technique.
-Firstly, while our method reduces the complexity of LLMs and enhances computational efficiency, there is an inevitable balance between sparsity and performance that requires careful calibration.
-Additionally, the effectiveness of AdaGP, similar to any pruning method, may vary across different models and tasks, and its generalizability to all scenarios remains an area for further exploration.
-Lastly, our approach assumes certain structural properties of the neural network, such as layer-wise decomposability, which may not hold for all architectures.
-These limitations, while not undermining the contributions of AdaGP, suggest avenues for future research and refinement.
-Acknowledgement
-This research used resources of the Argonne Leadership Computing Facility at Argonne National Laboratory, which is supported by the Office of Science of the U.S.
-Department of Energy under contract DE-AC02-06CH11357.
-We gratefully acknowledge the computing resources provided on Swing, a high-performance computing cluster operated by the Laboratory Computing Resource Center at Argonne National Laboratory.
-Government license
-The submitted manuscript has been created by UChicago Argonne, LLC, Operator of Argonne National Laboratory ("Argonne”).
-Argonne, a U.S.
-Department of Energy Office of Science laboratory, is operated under Contract No.
-DE-AC02-06CH11357.
-The U.S.
-Government retains for itself, and others acting on its behalf, a paid-up nonexclusive, irrevocable worldwide license in said article to reproduce, prepare derivative works, distribute copies to the public, and perform publicly and display publicly, by or on behalf of the Government.
-The Department of Energy will provide public access to these results of federally sponsored research in accordance with the DOE Public Access Plan (http://energy.gov/downloads/doe-public-access-plan)
+The versatility and minimal computational overhead of SparseLLM complement its integration with current pruning technologies, underscoring its potential as a universal tool for enhancing the performance and accessibility of LLMs.
 Appendix A Appendix
-A.1 OPT-125m
-Detailed results for OPT-125m are in Table 3.
-For a thorough investigation, we additionally consider unstructured sparsity of 50% and 60% as well as pruning the full model (instead of the first 50% of the transformer decoder layers).
-Similar to other models, we report the perplexity results for the un-pruned dense model next to the name of the model in the table.
-In particular, we see that SparseGPT and AdaGP outperform Magnitude and Wanda with a significant margin across different level of sparsity.
-AdaGP achieves similar perplexity with SparseGPT at smaller sparsity, but demonstrates much better perplexity for larger sparsity.
-Similar perplexity trends are observed across the three datasets, although, PTB, having the highest perplexity for each sparsity and method, seems to be the most challenging dataset among the three.
-OPT-125m (WikiText2: 27.66; PTB: 38.99; C4: 26.56)
-Sparsity Method WikiText2 PTB C4 Sparsity Method WikiText2 PTB C4
-50% Magnitude 193.39 276.33 141.03 80% Magnitude 4890.96 4121.49 3213.85
-Wanda 39.58 57.44 35.12 Wanda 1912.45 2512.93 1066.86
-SparseGPT 37.49 55.91 33.63 SparseGPT 2072.12 1952.85 1050.83
-AdaGP 37.09 55.82 33.40 AdaGP 1358.10 1418.09 654.54
-60% Magnitude 920.29 1146.49 598.45 90% Magnitude 6613.18 5380.80 4475.29
-Wanda 76.30 106.92 61.33 Wanda 4940.89 4337.27 3126.02
-SparseGPT 62.22 87.70 49.85 SparseGPT 6131.57 6963.27 2443.33
-AdaGP 59.28 84.36 47.78 AdaGP 5291.64 5067.41 2003.09
-70% Magnitude 3806.96 3429.35 2263.37 2:4 Magnitude 342.51 798.99 223.19
-Wanda 351.83 412.52 248.94 Wanda 83.81 113.01 65.78
-SparseGPT 239.26 265.83 156.33 SparseGPT 61.44 92.30 51.42
-AdaGP 208.46 255.75 137.72 AdaGP 56.55 81.46 46.14
-Table 3: Perplexity on pruned OPT-125m; the lower the perplexity, the better
-A.2 Pseudo-code of AdaGP
-Input: Pre-trained weight matrices $\mathbf{W}_{\ell}$ and $\mathbf{W}_{\ell+1}$, inputs $\boldsymbol{a}_{\ell-1}^{pre}$ and $\boldsymbol{a}_{\ell}^{pre}$, and outputs $\mathbf{z}_{\ell}^{pre}$ and $\mathbf{z}_{\ell+1}^{pre}$ for FFN’s up-scaling and down-scaling layers $\ell$ and $\ell+1$; target sparsity $\rho$; constraint weight hyperparameters $\alpha$, $\beta$.
-1AdaGP():
+This section includes supplemental materials (pseudo-code, additional experiments, and plots).
+A.1 Pseudo-code of SparseLLM
+1
+Input: An OPT decoder layer containing FFN and MHA modules.
+FFN’s up-scaling linear layer pre-trained weight matrix $\mathbf{W}_{\ell}$, FFN’s down-scaling linear layer pre-trained weight matrix $\mathbf{W}_{\ell+1}$, input of the up-scaling linear layer $\boldsymbol{a}_{\ell-1}^{pre}$, output of the down-scaling linear layer $\mathbf{z}_{\ell+1}^{pre}$, target sparsity $\rho$, constraint weight hyperparameters $\alpha$, $\beta$.
+2 SparseLLM on FFN():
+3
 4       Initialize $\mathbf{z}_{\ell}=\mathbf{z}_{\ell}^{pre}$, $\boldsymbol{a}_{\ell}=\boldsymbol{a}_{\ell}^{pre}$triangleright Initialize slack variables
@@
 6       for step $i=1,\cdots,K$ do
-5             $\mathbf{W}_{\ell}=\mathbf{z}_{\ell}\boldsymbol{a}_{\ell-1}^{\dagger}$, $\mathbf{W}_{\ell+1}=\mathbf{z}_{\ell+1}^{pre}\boldsymbol{a}_{\ell}^{\dagger}$
+7
+8             $\mathbf{W}_{\ell}=\mathbf{z}_{\ell}\boldsymbol{a}_{\ell-1}^{\dagger}$,
+9             $\mathbf{W}_{\ell+1}=\mathbf{z}_{\ell+1}\boldsymbol{a}_{\ell}^{\dagger}$
 10             $\mathbf{M}_{\ell},\widehat{\mathbf{W}}_{\ell}=\operatorname*{arg\,min}\|\mathbf{W}_{\ell}\boldsymbol{a}_{\ell-1}^{pre}-(\mathbf{M}_{\ell}\odot\widehat{\mathbf{W}}_{\ell})\boldsymbol{a}_{\ell-1}^{pre}\|_{2}^{2}$triangleright Prune layer $\ell$ by SparseGPT solver
 11             $\mathbf{M}_{\ell+1},\widehat{\mathbf{W}}_{\ell+1}=\operatorname*{arg\,min}\|\mathbf{W}_{\ell+1}\boldsymbol{a}_{\ell}-(\mathbf{M}_{\ell+1}\odot\widehat{\mathbf{W}}_{\ell+1})\boldsymbol{a}_{\ell}\|_{2}^{2}$triangleright Prune layer $\ell+1$ by SparseGPT solver
-8             $\mathbf{W}_{\ell+1}=\mathbf{M}_{\ell+1}\odot\widehat{\mathbf{W}}_{\ell+1}$, $\mathbf{W}_{\ell}=\mathbf{M}_{\ell}\odot\widehat{\mathbf{W}}_{\ell}$
+12             $\mathbf{W}_{\ell+1}=\mathbf{M}_{\ell+1}\odot\widehat{\mathbf{W}}_{\ell+1}$,
+13             $\mathbf{W}_{\ell}=\mathbf{M}_{\ell}\odot\widehat{\mathbf{W}}_{\ell}$
 14             $\boldsymbol{a}_{\ell}=(\alpha\mathbf{W}_{\ell+1}^{\intercal}\mathbf{W}_{\ell+1}+\beta\mathbf{I})^{-1}(\alpha\mathbf{W}_{\ell+1}^{\intercal}\mathbf{z}_{\ell+1}^{pre}+\beta\phi_{\ell}(\mathbf{z}_{\ell}))$triangleright Update activations
@@
 18                         ${(\mathbf{z}_{\ell})}_{j}={(\mathbf{z}_{\ell}^{(1)})}_{j}$triangleright Update outputs
-14
+19
 20                  else
 21                        ${(\mathbf{z}_{\ell})}_{j}={(\mathbf{z}_{\ell}^{(2)})}_{j}$triangleright Update outputs
-17
-18
+22
+23
+24
 25      return $\mathbf{W}_{\ell}$, $\mathbf{W}_{\ell+1}$
-Algorithm 1 Adaptive Global Pruning for the FFN module of an LLM.
-The Adaptive Global Pruning (AdaGP) algorithm presented in Algorithm 1 focuses on the optimization of the Feed-Forward Network (FFN) module within a Large Language Model (LLM).
-The key inputs to the algorithm include the pre-trained weight matrices, inputs and outputs for both the up-scaling and down-scaling linear layers of the FFN, along with a set of hyperparameters and a target sparsity.
-The goal of AdaGP is to achieve a targeted level of sparsity in the linear layers without significantly compromising the model’s performance.
-Initiating with the pre-trained weights, AdaGP employs a series of pruning and activation update steps across $K$ epochs.
-In each epoch, it solves optimization problems to prune the weights in both up-scaling and down-scaling layers, followed by updating the activation variables.
-The utilization of SparseGPT solver for pruning and the strategic update of activations ensure that the pruned network behaves as similarly as possible to that of the original network.
-The final output of the algorithm is a pair of pruned weight matrices for the consecutive layers, which are expected to deliver comparable, if not improved, performance with a reduced number of non-zero parameters.
-Figure 4: Visualization of the AdaGP Algorithm Applied to the FFN Module in a transformer decoder.
-This diagram illustrates the flow of data through the layers of the decoder, where AdaGP performs global pruning to enhance sparsity and computational efficiency without compromising the model’s performances.
-Each decoder layer is shown with its core components: Masked Multi-Head Attention and Feed Forward network, with AdaGP focusing on the latter for pruning.
-For other linear layers outside of the FFN modules, AdaGP follows the same local pruning as SparseGPT does to prune them.
-A.3 Visualization of AdaGP for FFN Global Pruning
-Figure 4 presents a schematic representation of the AdaGP algorithm applied to the Feed-Forward Network (FFN) module within a transformer model’s decoder layer.
-This visualization highlights the process of global pruning applied to the FFN module, aimed at reducing the model’s complexity while preserving its ability to capture essential patterns in the data.
-The AdaGP method operates within the decoder block, selectively pruning the weights of the FFN module.
-Each decoder block comprises a series of operations.
-It starts with masked multi-head attention and follows by additive normalization.
-Subsequently, it passes through the feed-forward network in which the pruning happens with AdaGP.
-The figure illustrates the flow of information from an input token through successive layers of the decoder.
-AdaGP not only enhances computational efficiency but also has the potential to improve model generalization by eliminating redundant parameters.
-A.4 Calibration Samples
-Figure 5 presents how perplexity changes with the calibration sample sizes on the datasets PTB and C4, respectively.
-On one hand, in both plots, as the number of calibration samples increases, the perplexity decreases for both SparseGPT and AdaGP.
-This suggests that having more calibration samples can be beneficial in the pruning process.
+26
+27 SparseLLM on MHA():
+28       for each linear layer $\ell$ in MHA module do
+29
+30             Fix $\mathbf{z}_{\ell}=\mathbf{z}_{\ell}^{pre}$, $\boldsymbol{a}_{\ell}=\boldsymbol{a}_{\ell}^{pre}$triangleright Fix intermediate variables
+31             $\mathbf{M}_{\ell},\widehat{\mathbf{W}}_{\ell}=\operatorname*{arg\,min}\|\mathbf{W}_{\ell}\boldsymbol{a}_{\ell-1}^{pre}-(\mathbf{M}_{\ell}\odot\widehat{\mathbf{W}}_{\ell})\boldsymbol{a}_{\ell-1}^{pre}\|_{2}^{2}$triangleright Prune layer $\ell$ by SparseGPT solver
+32
+33
+34      return $\mathbf{W}_{\ell}$} for all linear layers in MHA
+35
+Algorithm 1 SparseLLM Pruning of OPT Models.
+The SparseLLM algorithm presented in Algorithm 1 demonstrates how SparseLLM works on an OPT decoder layer.
+The key inputs to the algorithm include the pre-trained weight matrices for both the up-scaling and down-scaling linear layers of the FFN, along with a set of hyperparameters and a sparsity constraint.
+The goal of SparseLLM is to achieve a targeted level of sparsity in the linear layers without significantly compromising the model’s performance.
+Initiating with the pre-trained weights, SparseLLM employs a series of pruning and activation update steps across $K$ iterations.
+In each iteration, it solves optimization problems to prune the current and subsequent layer weights, followed by updating the activation variables.
+The utilization of SparseGPT solvers for pruning and the strategic update of activations ensures that the pruned network approximates the original network’s behavior as closely as possible.
+The final output of the algorithm is a pair of pruned weight matrices for the consecutive layers, which are expected to deliver comparable or improved performance with a reduced number of parameters.
+A.2 Two-layer Demo on the Details behind our Global Pruning
+Figure 4 illustrates the SparseLLM pruning method compared to conventional global pruning and local pruning, using a two-layer neural network as an abstraction for simplicity.
+The figure is divided into three main parts:
+On the left, conventional global pruning is depicted.
+This method applies a global mask to the entire network, resulting in significant memory costs due to poor scalability.
+Both functions $f_{1}$ and $f_{2}$ are pruned using the same mask across all layers, leading to high memory usage.
+In the middle, local pruning is shown, where each layer is pruned independently.
+This approach reduces memory costs by applying separate masks to each layer.
+However, it inevitably sacrifices performance because it ignores global supervision, which can lead to suboptimal pruning decisions that do not consider the network as a whole.
+On the right, the adaptive global pruning method of SparseLLM is presented.
+This method achieves global pruning with low memory cost by leveraging auxiliary variables and soft constraints.
+It combines the benefits of global pruning—considering the entire network structure—with efficient memory usage.
+The introduction of auxiliary variables allows for flexible and adaptive pruning, ensuring that the overall performance of the network is maintained while keeping memory costs low.
+Thus, the figure highlights the trade-offs between different pruning strategies.
+Conventional global pruning incurs high memory costs, local pruning reduces memory usage at the expense of performance, and the adaptive global pruning of SparseLLM strikes a balance by maintaining performance with lower memory requirements through the use of auxiliary variables and soft constraints.
+Figure 4: Illustration of SparseLLM pruning method compared to conventional global pruning and local pruning.
+We consider a two-layer neural network as an abstraction for simplicity.
+Global pruning (left) is memory prohibitive due to poor scalability.
+Local pruning (mid) considers pruning each layer independently, while inevitably sacrificing performance due to the ignorance of global supervision.
+Our adaptive global pruning (right) achieves global pruning with low memory cost by leveraging auxiliary variables and soft constraints.
+A.3 Calibration Samples
+Figure 5 and Figure 6 present how perplexity changes with the calibration sample sizes on the datasets PTB and C4 for OPT-2.7b and LlaMA-2 7B, respectively.
+In both figures, as the number of calibration samples increases, the perplexity decreases for both SparseGPT and SparseLLM.
+This indicates that having more calibration samples can be beneficial in the pruning process.
 Perplexity decreases more rapidly from 8 samples to 32 samples.
 Beyond 32 samples, the rate at which perplexity decreases starts to slow down.
-For PTB dataset, the perplexity even increases from 64 samples to 128 samples.
-On the other hand, increasing the number of calibration samples requires more computational resources, e.g., memory and computation time, in the execution.
-A calibration sample size should be between 32 and 64 can therefore strike a balance between good performances and computational efficiency.
-Lastly, we want to point out that the figures show that AdaGP starts to achieve better perplexity than SparseGPT does from 32 samples and remains better for 64 and 128 samples.
+In addition, increasing the number of calibration samples requires more computational resources, e.g., memory and computation time, in the overall pruning process.
+This suggests that the calibration sample sizes should be between 32 and 64 to ensure good performance while maintaining computational efficiency.
+Lastly, the figures show that SparseLLM achieves better perplexity than SparseGPT does with 32 or larger sample sizes for both OPT and LlaMA-2 models.
 Figure 5: Sensitivity of OPT-2.7b on the calibration sample sizes for datasets PTB and C4.
-A.5 Hyperparameter $\alpha$ and $\beta$ Selection
-Hyperparameters $\alpha$ and $\beta$ are used in Eq.
-5.
-We select $\alpha$ and $\beta$ from the set $\{0.1,1.0,5.0\}$ and perform a study on model OPT-2.7b to understand the impact of the hyperparameters.
-We determine that $\alpha=\beta=0.1$ give the best and numerically stable performances.
+Figure 6: Sensitivity of LlaMA-2 7b models on the calibration sample sizes for datasets PTB and C4.
+A.4 Computation Time vs. Model Sizes
+We study how the computation time of SparseGPT and SparseLLM varies with different model sizes, as illustrated in Table 3 and Table 4 for OPT models and LlaMA-2 models.
+The rate at which the time taken increases is comparable for SparseGPT and SparseLLM as the model size increases.
+Additionally, computation time for SparseLLM are reported for a configuration of 4 to 10 epochs.
+As we have reported in Section 5, SparseLLM can reduce the training loss in as few as 2 to 3 epochs.
+This suggests that the proposed SparseLLM remains computationally efficient.
+Table 3: Computation time in seconds of OPT models.
+Method OPT-125m OPT-1.3b OPT-2.7b OPT-6.7b OPT-13b OPT-30b OPT-66b
+SparseGPT 1.47 4.03 4.36 8.15 10.46 20.94 34.78
+SparseLLM 19.58 74.76 97.44 184.96 209.26 164.34 270.38
+Table 4: Computation time in seconds of LlaMA-2 models.
+Method Llama-2 7b Llama-2 13b
+SparseGPT 179.08 314.95
+SparseLLM 2201.99 4797.07
+A.5 Experiment Results for Additional Models
+Detailed results on perplexity and zero-shot task accuracy for additional models are reported in Table 5 and Table 6.
+Similar to other models, we report the perplexity results for the dense model next to the name of the model in the table.
+In particular, we see that SparseGPT and SparseLLM outperform Magnitude and Wanda with a significant margin across different sparsity.
+SparseLLM shares similar perplexity with SparseGPT for smaller sparsity but demonstrates much better perplexity for larger sparsity.
+Similar perplexity trends are observed across all three datasets, although, PTB, having the highest perplexity for each sparsity and method, is likely the most challenging dataset among the three.
+For the zero-shot taks accuracy, we see that SparseLLM achieves comparable results to SparseGPT for smaller perplexity and the performance improvements are more obvious and significant with higher sparsity.
+Table 5: Perplexity in high sparsity regimes ($\geq 70\%$); the lower the perplexity, the better.
+OPT-125m (WikiText2 (WT2): 27.66; PTB: 38.99; C4: 26.56)
+Sparsity 70% 80% 90% 3:4
+Dataset WT2 PTB C4 WT2 PTB C4 WT2 PTB C4 WT2 PTB C4
+Magnitude 3806.96 3429.35 2263.37 4890.96 4121.49 3213.85 6613.18 5380.80 4475.29 - - -
+Wanda 351.83 412.52 248.94 1912.45 2512.93 1066.86 4940.89 4337.27 3126.02 - - -
+SparseGPT 239.26 265.83 156.33 2072.12 1952.85 1050.83 6131.57 6963.27 2443.33 1482.61 2215.44 657.26
+SparseLLM 208.46 255.75 137.72 1358.10 1418.09 654.54 5291.64 5067.41 2003.09 914.87 1210.84 450.01
+OPT-6.7b (WikiText2 (WT2): 10.86; PTB: 15.77; C4: 12.71)
+Sparsity 70% 80% 90% 3:4
+Dataset WT2 PTB C4 WT2 PTB C4 WT2 PTB C4 WT2 PTB C4
+Magnitude 7054.21 5437.44 4850.25 7937.49 5971.86 6031.54 2.4e4 2.5e4 2.1e4 - - -
+Wanda 54.95 129.73 116.67 1493.58 1196.93 996.00 2.1e4 2.0e4 1.8e4 - - -
+SparseGPT 12.27 18.90 15.28 31.04 51.26 29.42 8871.24 5713.57 3797.20 570.08 361.81 328.18
+SparseLLM 12.16 18.39 14.93 23.96 39.32 26.97 2095.85 1842.48 953.44 83.36 128.99 62.11
+OPT-13b (WikiText2 (WT2): 10.13; PTB: 14.52; C4: 12.06)
+Sparsity 70% 80% 85% 3:4
+Dataset WT2 PTB C4 WT2 PTB C4 WT2 PTB C4 WT2 PTB C4
+Magnitude 9037.12 7734.58 5909.47 1.1e4 9140.88 6340.22 1.3e4 1.3e4 9087.50 - - -
+Wanda 30.94 39.26 33.31 4216.04 2894.77 2450.57 1.1e4 1.1e4 7244.96 - - -
+SparseGPT 10.89 16.35 13.39 21.42 33.62 21.01 8408.03 6380.30 3416.23 4715.16 7454.37 2.11e4
+SparseLLM 10.96 16.57 13.38 19.07 28.77 19.29 2052.27 1536.51 538.61 289.17 687.48 677.13
+Table 6: Accuracy (%) of zero-shot tasks; the higher the accuracy, the better.
+OPT-6.7b
+Sparsity Method BoolQ RTE HellaSwag WinoGrande ARC-e ARC-c OBQA Mean
+Dense 66.12 56.03 50.49 65.27 65.72 30.63 27.60 51.69
+70% SparseGPT 61.74 54.87 48.46 63.85 64.31 29.27 25.40 49.70
+SparseLLM 60.61 54.51 48.8 62.9 64.14 30.03 26.60 49.66
+80% SparseGPT 55.08 48.38 42.22 59.43 57.79 25.85 21.40 44.31
+SparseLLM 58.69 51.26 43.78 59.67 58.38 26.88 22.00 45.81
+90% SparseGPT 38.53 53.07 26.00 48.07 26.81 21.67 14.40 32.65
+SparseLLM 46.48 52.71 26.21 51.70 27.44 19.71 13.40 33.95
+3:4 SparseGPT 46.70 54.15 28.82 51.07 32.45 18.17 15.40 35.25
+SparseLLM 53.49 53.42 36.24 53.51 43.94 22.61 17.40 40.09
+OPT-13b
+Sparsity Method BoolQ RTE HellaSwag WinoGrande ARC-e ARC-c OBQA Mean
+Dense 65.87 57.76 52.44 66.02 67.82 33.46 28.62 53.14
+70% SparseGPT 63.03 54.87 50.89 65.43 67.47 32.85 26.40 51.56
+SparseLLM 63.85 55.23 50.73 65.67 66.46 31.83 27.20 51.57
+80% SparseGPT 59.72 52.35 46.82 61.48 62.50 31.23 21.80 47.99
+SparseLLM 60.89 53.07 46.19 62.12 62.21 30.38 23.00 48.27
+90% SparseGPT 47.49 52.71 33.17 51.54 39.98 21.33 17.80 37.72
+SparseLLM 53.43 52.71 38.19 52.96 46.68 25.26 17.40 40.95
+3:4 SparseGPT 47.55 53.43 31.30 50.20 37.63 22.53 17.60 37.18
+SparseLLM 51.13 52.35 38.51 55.96 49.24 24.83 21.40 41.92

### v4

@@
 Addressing these challenges, we propose SparseLLM, a novel framework that redefines the global pruning process into manageable, coordinated subproblems, allowing for resource-efficient optimization with global optimality.
-SparseLLM’s approach, which conceptualizes LLMs as a chain of modular functions and leverages auxiliary variables for problem decomposition, not only facilitates a pragmatic application on LLMs but also demonstrates significant performance improvements, particularly in high-sparsity regimes where it surpasses current state-of-the-art methods.
-Code can be found at https://github.com/BaiTheBest/SparseLLM.
+SparseLLM’s approach, which conceptualizes LLMs as a chain of modular functions and leverages auxiliary variables for problem decomposition, not only facilitates a pragmatic application on LLMs but also demonstrates significant performance improvements, particularly in high-sparsity regimes, surpassing current state-of-the-art methods.
+Our source code is publicly available at https://github.com/BaiTheBest/SparseLLM.
 1 Introduction
@@
 To make LLMs more accessible, extensive efforts have been devoted to model compression of LLMs Xu and McAuley (2023); Bai et al.
-(2024), including pruning, quantization, knowledge distillation, and low-rank factorization.
+(2024a), including pruning, quantization, knowledge distillation, and low-rank factorization.
 Pruning, by introducing sparsity, jointly enhances memory and computational efficiency and offers unparalleled flexibility, seamlessly integrating with any LLMs, thus standing out as a highly effective and widely adopted compression strategy.
@@
 However, conventional pruning techniques, which rely on global pruning and require loading the entire model into the same GPU Mallya and Lazebnik (2018); Singh and Alistarh (2020), become impractical for today’s LLMs due to their vast size.
-Recently, several local pruning methods, which consider compressing each layer separately, and the overall compressed model is then obtained by “stitching together” the individually compressed layers, are proposed for billion-scale LLMs.
-SparseGPT Frantar and Alistarh (2023), an efficient unstructured pruning method for LLMs with hundreds of billions of parameters, achieves up to 60% parameter reduction with minimal performance loss.
+Recently, several local pruning methods have been proposed for billion-scale LLMs.
+These methods compress each layer separately, and the overall compressed model is then obtained by “stitching together” the individually compressed layers.
+SparseGPT Frantar and Alistarh (2023), an efficient unstructured pruning method for LLMs with hundreds of billions of parameters, achieves parameter reduction of up to 60% with minimal performance loss.
 Another approach, Wanda Sun et al.
 (2023), introduces a novel pruning criterion that evaluates weights by considering both magnitude and related input activations.
-Despite its efficiency gains, local pruning only aims to minimize the local error for each specific layer under sparsity constraints, resulting in a suboptimal solultion for the overall model.
+Despite its efficiency gains, local pruning only aims to minimize the local error for each specific layer under sparsity constraints, resulting in a suboptimal solution for the overall model.
 This is because local pruning over-aligns the intermediate layers’ activations, leading to suboptimal performance, especially in high-sparsity regimes Singh and Alistarh (2020); Sung et al.
@@
 LLM pruning techniques can broadly be categorized into structured and unstructured prunings.
-Unstructured pruning looks at simplifying the complexity of LLMs by removing certain parameters regardless of the model’s inherent structure.
+Unstructured pruning Zhang et al.
+(2024); Bai et al.
+(2024b) looks at simplifying the complexity of LLMs by removing certain parameters regardless of the model’s inherent structure.
 This approach typically involves setting a threshold to nullify parameters below it, leading to a model with a non-uniform sparse structure.
@@
 ZipLM Kurtic et al.
-(2023), a structured compression method for LLMs, is proposed to optimize for compression and accuracy while considering specific hardware constraints.
+(2023), a structured pruning method for LLMs, is proposed to optimize for compression and accuracy while considering specific hardware constraints.
+More recently, Xia et al introduced LLM-shearing Xia et al.
+(2023), a structured pruning method that scales down LLaMA models by selectively pruning layers, heads, and dimensions.
+This approach, combined with dynamic data batching, reduces pre-training compute costs while maintaining competitive performance, outperforming similar open-source models on key tasks.
 Our work falls in the category of unstructured pruning of LLMs, where existing methods such as SparseGPT and Wanda only consider an entirely local pruning algorithm and suffer from suboptimal performance.
@@
 This strategy strikes a balance between the computational feasibility of pruning large-scale models and the effectiveness of the pruning process, adhering to the limitations and practices of state-of-the-art LLM pruning frameworks.
-Formally speaking, rather than try to solve Eq. 3 directly, we first relax the constraints by adding an $\ell_{2}$-penalty function to the objective and attack the unconstrained problem:
+Formally speaking, rather than trying to solve Eq. 3 directly, we first relax the constraints by adding an $\ell_{2}$-penalty function to the objective and attack the unconstrained problem:
 $\begin{split}\mathcal{L}(\mathbf{z}_{L},\mathbf{y})+\alpha\sum\nolimits_{\ell\in[L]}\|\mathbf{z}_{\ell}-(\mathbf{M}_{\ell}\odot\widehat{\mathbf{W}}_{\ell})\boldsymbol{a}_{\ell-1}\|_{2}^{2}+\beta\sum\nolimits_{\ell\in\Omega_{\text{FFN}}}\|\boldsymbol{a}_{\ell}-\phi_{\ell}(\mathbf{z}_{\ell})\|_{2}^{2},\end{split}$ (4)
@@
 The careful and elaborate design of our Eq.
-5 allows us to derive a closed-form solution to every subproblem as shown in the following.
+5 allows us to derive a closed-form solution to every subproblem as shown below.
 Pruning weight.
@@
 5, under the condition that the activation function $\phi$ is ReLU.
-Notice that (1) the objective function is convex with respect to each variable when all others are fixed, and (2) given that closed-form solutions exist for the subproblems in the alternating optimization scheme, the proposed algorithm is guaranteed to converge to a global optimum.
+Notice that (1) the objective function is convex with respect to each variable when all others are fixed, and (2) given that closed-form solutions exist for the subproblems in the alternating optimization scheme, the proposed algorithm resembles multiblock ADMM which has been shown to converge to in many applications.
 4.3.3 Pruning of MHAs
@@
 Using SparseGPT as the solver, the exact pruning step has a complexity of $O(h^{3})$.
-The second step, updating activation, involves matrix inversion of the weight matrix $\mathbf{W}_{\ell}$ (size $h\times h$) with a complexity of $O(h^{3})$.
+The second step, updating activations, involves matrix inversion of the weight matrix $\mathbf{W}_{\ell}$ (size $h\times h$) with a complexity of $O(h^{3})$.
 The third step, updating outputs, has a lower complexity.
-Thus, the overall algorithm complexity is bounded by $O(h^{3})$, making our method’s per-epoch time complexity comparable to SparseGPT.
+Thus, the overall algorithm complexity is bounded by $O(h^{3})$, therefore making our method’s per-epoch time complexity comparable to SparseGPT.
 5 Experiments
 Experiments setup.
-We implement SparseLLM in PyTorch Paszke et al.
+We implemented SparseLLM in PyTorch Paszke et al.
 (2019) and use the HuggingFace Transformers library Wolf et al.
@@
 (2020) dataset.
-This represents generic text data crawled from the internet and ensures our experiments remain zero-shot as no task-specific data is seen during pruning.
+This represents generic text data crawled from the internet and ensures our experiments are zero-shot as no task-specific data is seen during pruning.
 We followed existing work Frantar and Alistarh (2023); Sun et al.
@@
 (2022) and LlaMA-2 model family Touvron et al.
-(2023) in our experiments.
+(2023) in our experiments as well as the most recent LlaMA-3 model.
 We show results on different sizes of models to provide a broader picture for the performances of SparseLLM.
@@
 (2023).
-Table 1: Perplexity in high sparsity regimes ($\geq 70\%$); the lower the perplexity, the better.
+Comparison methods.
+We compare against three baselines, magnitude pruning Zhu and Gupta (2017) applied locally, and two other state-of-the-art local pruning methods, SparseGPT Frantar and Alistarh (2023) and Wanda Sun et al.
+(2023).
+5.1 Results and analyses
+Pruning vs.
+model sizes.
+We begin by exploring the pruning capabilities of SparseLLM across various model sizes in comparison to baseline methods.
+For each model, we consider unstructured sparsity ranging from 70% to 90% with a 10% increment, as well as a 3:4 semi-structured sparsity.
+The 3:4 semi-structured sparsity is inspired by our preliminary results that suggest good performance SparseLLM at high sparsity regimes.
+However, note that two of our baselines, Magnitude and Wanda, are unable to be configured to this sparsity out-of-box.
+We conduct a sensitivity study on the calibration sample sizes (see Appendix A.3) and use calibration sample sizes between 32 and 64 for all experiments.
+Moreover, we prune the first 50% of the Transformer decoder layers in each model to achieve a balance between the computation resources and the performances.
+Detailed results can be found in Table 1 and Table 2 as well as Table 8 in Appendix A.5.
+Note that in Table 2 for LlaMA-3 model, we only compare SparseGPT to the proposed SparseLLM.
+The perplexity results of the dense models are reported next to the names of the models.
+Table 1: Perplexity of OPT models for sparsity $\geq 70\%$; the lower the perplexity, the better.
 OPT-1.3b (WikiText2 (WT2): 14.62; PTB: 20.29; C4: 16.07)
@@
 SparseLLM 13.82 21.07 16.14 23.87 37.09 24.90 1200.12 759.11 527.70 56.90 77.14 52.77
+OPT-13b (WikiText2 (WT2): 10.13; PTB: 14.52; C4: 12.06)
+Sparsity 70% 80% 85% 3:4
+Dataset WT2 PTB C4 WT2 PTB C4 WT2 PTB C4 WT2 PTB C4
+Magnitude 9037.12 7734.58 5909.47 1.1e4 9140.88 6340.22 1.3e4 1.3e4 9087.50 - - -
+Wanda 30.94 39.26 33.31 4216.04 2894.77 2450.57 1.1e4 1.1e4 7244.96 - - -
+SparseGPT 10.89 16.35 13.39 21.42 33.62 21.01 8408.03 6380.30 3416.23 4715.16 7454.37 2.11e4
+SparseLLM 10.96 16.57 13.38 19.07 28.77 19.29 2052.27 1536.51 538.61 289.17 687.48 677.13
 OPT-30b (WikiText2 (WT2): 9.56; PTB: 14.04; C4: 11.45)
@@
 SparseLLM 9.37 13.66 11.37 16.45 21.00 17.70 7504.17 5644.65 3683.91 4641.8 5296.93 1618.43
+Table 2: Perplexity of LlaMA models for sparsity $\geq 70\%$; the lower the perplexity, the better.
 LlaMA-2 7b (WikiText2 (WT2): 5.47; PTB: 37.91; C4: 7.26)
@@
 SparseLLM 12.95 277.76 15.77 36.36 578.35 38.63 646.15 1078.94 466.98 53.71 632.11 50.40
-Table 2: Accuracy (%) of zero-shot tasks; the higher the accuracy, the better.
+LlaMA-3 8b (WikiText2 (WT2): 6.14; PTB: 11.18; C4: 9.45)
+Sparsity 70% 80% 90% 3:4
+Dataset WT2 PTB C4 WT2 PTB C4 WT2 PTB C4 WT2 PTB C4
+SparseGPT 22.37 36.56 30.53 72.87 113.95 79.86 214.68 261.18 198.34 96.75 107.52 102.11
+SparseLLM 20.98 33.78 28.94 57.83 85.98 72.18 197.47 241.68 181.69 76.33 99.54 93.68
+Table 3: Perplexity of 2:4 sparsity; the lower the perplexity, the better.
+OPT-1.3b OPT-2.7b OPT-6.7b OPT-13b
+Dataset WT2 PTB C4 WT2 PTB C4 WT2 PTB C4 WT2 PTB C4
+Magnitude 96.68 133.92 48.08 272.34 308.55 267.70 64.11 92.23 82.67 67.07 110.77 52.61
+Wanda 15.63 24.04 18.23 13.66 21.67 16.10 11.86 18.54 14.77 10.33 15.35 12.54
+SparseGPT 15.11 23.71 17.88 12.62 19.28 15.12 11.30 16.90 13.51 10.20 15.14 12.48
+SparseLLM 14.97 23.40 17.67 12.62 19.28 15.12 11.07 16.73 13.42 10.20 15.14 12.41
+From the tables, it shows a general trend of increasing perplexity with increasing sparsity.
+Moreover, we observe a trend of decreasing perplexity for SparseGPT and SparseLLM at the same sparsity with increasing model sizes.
+However, such a trend is not obvious for Magnitude and Wanda.
+We also observe that SparseGPT and SparseLLM consistently outperform Magnitude and Wanda by a significant margin.
+For smaller sparsity, SparseLLM achieves comparable perplexity to SparseGPT.
+As we increase the sparsity, SparseLLM starts to demonstrate noticeable improvements over SparseGPT.
+In numerous instances for the OPT model family, SparseLLM achieves perplexity reductions of more than 50% compared to SparseGPT.
+We also see that performance improvements from SparseLLM over SparseGPT are more significant for the OPT model family than the LlaMA-2 model family.
+We provide additional set of perplexity results for a 2:4 semi-structured sparsity for a few OPT models in Table 3.
+We see that SparseLLM and SparseGPT generally outperform Magnitude and Wanda while SparseLLM has comparable if not slightly better performances compared to SparseGPT with the 2:4 semi-structured sparsity.
+Note that a 2:4 semi-structure sparsity is considered to be in low sparsity regime.
+Figure 3: Fast convergence of SparseLLM. Training loss per epoch for pruning layer 3 of OPT-125m at 80% sparsity (Left) and layer 6 of LlaMA-2 13b at 70% sparsity (Right).
+Zero-shot experiments.
+To further conclude the evaluations and discussions, we show results for several zero-shot tasks in Table 4 and Table 5 as well as Table 9 in Appendix A.5, comparing SparseGPT and SparseLLM.
+These evaluations are known to be relatively noisy Dettmers et al.
+(2022), but more interpretable.
+We also report the results for zero-shot tasks from the dense models in the “Dense" row.
+We see that the accuracy of both methods decreases with increasing sparsity, which is expected, as more parameters are pruned.
+A similar trend of increasing accuracy with increasing model size is observed too.
+Across all the tasks, OBQA and ARC-c remain the most challenging ones as the accuracy for both methods is 30% or below 30% while both methods perform well for BoolQ, RTE, WinoGrande, and ARC-e.
+In general, SparseLLM is able to achieve higher accuracy in the majority of tasks across the models of different sizes in both OPT and LlaMA-2 model families.
+Table 4: Accuracy (%) of zero-shot tasks for OPT models; the higher the accuracy, the better.
+OPT-13b
+Sparsity Method BoolQ RTE HellaSwag WinoGrande ARC-e ARC-c OBQA Mean
+Dense 65.87 57.76 52.44 66.02 67.82 33.46 28.62 53.14
+70% SparseGPT 63.03 54.87 50.89 65.43 67.47 32.85 26.40 51.56
+SparseLLM 63.85 55.23 50.73 65.67 66.46 31.83 27.20 51.57
+80% SparseGPT 59.72 52.35 46.82 61.48 62.50 31.23 21.80 47.99
+SparseLLM 60.89 53.07 46.19 62.12 62.21 30.38 23.00 48.27
+90% SparseGPT 47.49 52.71 33.17 51.54 39.98 21.33 17.80 37.72
+SparseLLM 53.43 52.71 38.19 52.96 46.68 25.26 17.40 40.95
+3:4 SparseGPT 47.55 53.43 31.30 50.20 37.63 22.53 17.60 37.18
+SparseLLM 51.13 52.35 38.51 55.96 49.24 24.83 21.40 41.92
 OPT-30b
@@
 SparseLLM 60.83 54.15 39.35 55.41 45.24 24.06 22.20 43.03
+Table 5: Accuracy (%) of zero-shot tasks for LlaMA models; the higher the accuracy, the better.
 LlaMA-2 7b
@@
 SparseLLM 61.71 55.71 28.56 51.62 32.11 18.49 13.8 37.43
-Comparison methods.
-We compare against three baselines, magnitude pruning Zhu and Gupta (2017) applied locally, and two other state-of-the-art local pruning methods, SparseGPT Frantar and Alistarh (2023) and Wanda Sun et al.
-(2023).
-5.1 Results and analyses
-Pruning vs.
-model sizes.
-We begin by exploring the pruning capabilities of SparseLLM across various model sizes in comparison to baseline methods.
-For each model, we consider unstructured sparsity ranging from 70% to 90% with a 10% increment, as well as a 3:4 semi-structured sparsity.
-The 3:4 semi-structured sparsity is inspired by our preliminary results that suggest good performance SparseLLM at high sparsity regimes.
-However, note that two of our baselines, Magnitude and Wanda, are unable to be configured to this sparsity out-of-box.
-We conduct a sensitivity study on the calibration sample sizes (see Appendix A.3) and use calibration sample sizes between 32 and 64 for all experiments.
-Moreover, we prune the first 50% of the Transformer decoder layers in each model to achieve a balance between the computation resources and the performances.
-Detailed results can be found in Table 1 and Table 5 (see Appendix A.5).
-The perplexity results of the dense models are reported next to the names of the models.
-Figure 3: Fast convergence of SparseLLM. Training loss per epoch for pruning layer 3 of OPT-125m at 80% sparsity (Left) and layer 6 of LlaMA-2 13b at 70% sparsity (Right).
-From the table, it shows a general trend of increasing perplexity with increasing sparsity.
-Moreover, we observe a trend of decreasing perplexity for SparseGPT and SparseLLM at the same sparsity with increasing model sizes.
-However, such a trend is not obvious for Magnitude and Wanda.
-We also observe that SparseGPT and SparseLLM consistently outperform Magnitude and Wanda by a significant margin.
-For smaller sparsity, SparseLLM achieves comparable perplexity to SparseGPT.
-As we increase the sparsity, SparseLLM starts to demonstrate noticeable improvements over SparseGPT.
-In numerous instances for the OPT model family, SparseLLM achieves perplexity reductions of more than 50% compared to SparseGPT.
-We also see that performance improvements from SparseLLM over SparseGPT are more significant for the OPT model family than the LlaMA-2 model family.
-Zero-shot experiments.
-To further conclude the evaluations and discussions, we show results for several zero-shot tasks in Table 2 and Table 6 (see Appendix A.5) comparing SparseGPT and SparseLLM.
-These evaluations are known to be relatively noisy Dettmers et al.
-(2022), but more interpretable.
-We also report the results for zero-shot tasks from the dense models in the “Dense" row.
-We see that the accuracy of both methods decreases with increasing sparsity, which is expected, as more parameters are pruned.
-A similar trend of increasing accuracy with increasing model size is observed too.
-Across all the tasks, OBQA and ARC-c remain the most challenging ones as the accuracy for both methods is 30% or below 30% while both methods perform well for BoolQ, RTE, WinoGrande, and ARC-e.
-In general, SparseLLM is able to achieve higher accuracy in the majority of tasks across the models of different sizes in both OPT and LlaMA-2 model families.
 Training loss vs.
@@
 By adeptly circumventing the scalability issues of global pruning and optimizing the local suboptimality of existing methods, SparseLLM stands as a significant advancement in the field.
-Our empirical results affirm its efficacy, particularly in high-sparsity environments, where it achieves a notable reduction in perplexity, thereby setting a new precedent for model compression.
+Our empirical results affirm its efficacy, particularly in high-sparsity environments.
+It achieves a notable reduction in perplexity, thereby setting a new precedent for model compression.
 The versatility and minimal computational overhead of SparseLLM complement its integration with current pruning technologies, underscoring its potential as a universal tool for enhancing the performance and accessibility of LLMs.
+Acknowledgments and Disclosure of Funding
+This work was supported by the National Science Foundation (NSF) Grant No.1755850, No.1841520, No.1942594, No.2403312, No.2007716, No.2007976, No.1907805.
+This work was supported by the U.S.
+Department of Energy, Office of Science, Advanced Scientific Computing Research, under Contract DE-AC02-06CH11357.
+This research used resources of the Argonne Leadership Computing Facility at Argonne National Laboratory, which is supported by the Office of Science of the U.S.
+Department of Energy under contract DE-AC02-06CH11357.
 Appendix A Appendix
@@
 A.4 Computation Time vs. Model Sizes
-We study how the computation time of SparseGPT and SparseLLM varies with different model sizes, as illustrated in Table 3 and Table 4 for OPT models and LlaMA-2 models.
+We study how the computation time per layer of SparseGPT and SparseLLM varies with different model sizes, as illustrated in Table 6 and Table 7 for OPT models and LlaMA-2 models.
 The rate at which the time taken increases is comparable for SparseGPT and SparseLLM as the model size increases.
@@
 Method OPT-125m OPT-1.3b OPT-2.7b OPT-6.7b OPT-13b OPT-30b OPT-66b
-SparseGPT 1.47 4.03 4.36 8.15 10.46 20.94 34.78
-SparseLLM 19.58 74.76 97.44 184.96 209.26 164.34 270.38
+SparseGPT 2.30 10.18 18.35 24.40 28.65 48.91 103.19
+SparseLLM 16.34 22.79 42.86 174.08 85.62 174.07 284.59
 Table 7: Computation time in seconds of LlaMA-2 models.
 Method Llama-2 7b Llama-2 13b
-SparseGPT 179.08 314.95
-SparseLLM 2201.99 4797.07
+SparseGPT 11.94 16.58
+SparseLLM 146.80 252.48
 A.5 Experiment Results for Additional Models
@@
 SparseLLM 12.16 18.39 14.93 23.96 39.32 26.97 2095.85 1842.48 953.44 83.36 128.99 62.11
-OPT-13b (WikiText2 (WT2): 10.13; PTB: 14.52; C4: 12.06)
-Sparsity 70% 80% 85% 3:4
-Dataset WT2 PTB C4 WT2 PTB C4 WT2 PTB C4 WT2 PTB C4
-Magnitude 9037.12 7734.58 5909.47 1.1e4 9140.88 6340.22 1.3e4 1.3e4 9087.50 - - -
-Wanda 30.94 39.26 33.31 4216.04 2894.77 2450.57 1.1e4 1.1e4 7244.96 - - -
-SparseGPT 10.89 16.35 13.39 21.42 33.62 21.01 8408.03 6380.30 3416.23 4715.16 7454.37 2.11e4
-SparseLLM 10.96 16.57 13.38 19.07 28.77 19.29 2052.27 1536.51 538.61 289.17 687.48 677.13
 Table 9: Accuracy (%) of zero-shot tasks; the higher the accuracy, the better.
@@
 SparseLLM 53.49 53.42 36.24 53.51 43.94 22.61 17.40 40.09
-OPT-13b
-Sparsity Method BoolQ RTE HellaSwag WinoGrande ARC-e ARC-c OBQA Mean
-Dense 65.87 57.76 52.44 66.02 67.82 33.46 28.62 53.14
-70% SparseGPT 63.03 54.87 50.89 65.43 67.47 32.85 26.40 51.56
-SparseLLM 63.85 55.23 50.73 65.67 66.46 31.83 27.20 51.57
-80% SparseGPT 59.72 52.35 46.82 61.48 62.50 31.23 21.80 47.99
-SparseLLM 60.89 53.07 46.19 62.12 62.21 30.38 23.00 48.27
-90% SparseGPT 47.49 52.71 33.17 51.54 39.98 21.33 17.80 37.72
-SparseLLM 53.43 52.71 38.19 52.96 46.68 25.26 17.40 40.95
-3:4 SparseGPT 47.55 53.43 31.30 50.20 37.63 22.53 17.60 37.18
-SparseLLM 51.13 52.35 38.51 55.96 49.24 24.83 21.40 41.92
+A.6 Hyperparameter Selection
+Hyperparameters $\alpha$ and $\beta$ are used in Eq.
+5.
+We select $\alpha$ and $\beta$ from the set $\{0.01,0.1,1,5,10,100\}$ and perform a study on models to understand the impact of the hyperparameters.
+Results for OPT-1.3b with 70% sparsity are shown in Table 10.
+Table 10: Ablations of the hyperparameters $\alpha$ and $\beta$ on OPT-1.3b with 70% sparsity (in perplexity)
+$\alpha$ / $\beta$ $0.01$ $0.1$ $1$ $5$ $10$ $100$
+$0.01$ 18.01 17.97 17.97 - - -
+$0.1$ 18.04 17.82 17.96 18.04 18.40 -
+$1$ 18.20 18.02 18.11 17.87 17.96 18.22
+$5$ 18.06 18.02 18.03 17.92 17.96 18.04
+$10$ 18.03 18.01 17.96 17.96 17.96 18.03
+$100$ 18.04 18.04 17.98 18.01 18.01 18.03
+A.7 Limitations and Future Work
+While SparseLLM marks a significant step forward in the efficient pruning of large language models, it is important to acknowledge the inherent trade-offs associated with any model compression technique.
+Firstly, while our method reduces the complexity of LLMs and enhances computational efficiency, there is an inevitable balance between sparsity and performance that requires careful calibration.
+Additionally, in this work, we still assume homogeneous sparsity, i.e., the pruning sparsity for each layer is the same and equal to the global sparsity.
+How to achieve heterogeneous sparsity under our framework and fully fulfill the potential of global pruning is of great interest.
+Lastly, the effectiveness of SparseLLM, like any pruning method, may vary across different models and tasks, and its generalizability to all scenarios remains an area for further exploration.